### v2

# Schrödinger operators with concentric $\delta$–shell interactions

Abstract
We study Schrödinger operators on $\mathbb{R}^{3}$ with finitely many concentric spherical $\delta$–shell interactions. The operators are defined via quadratic forms and characterized by continuity across each shell together with the standard jump condition for the normal derivative. Using a boundary integral approach based on the free Green kernel and single–layer potentials, we derive an explicit resolvent formula for an arbitrary number of shells with bounded coupling strengths. This yields a concrete Kreĭn–type representation and a boundary operator whose noninvertibility characterizes the discrete spectrum, and it is compatible with partial–wave reduction under rotational symmetry.
We then specialize to the two–shell case with constant couplings and obtain a detailed description of the negative spectrum. In particular, we prove that the ground state, when it exists, lies in the ss–wave sector and derive an explicit secular equation for bound states. For large shell separation, each bound level approaches the corresponding single–shell level with exponentially small corrections, while a genuine tunneling splitting appears when the single–shell levels are tuned to coincide. As a simple calibration, we relate the two–shell parameters to representative core–shell quantum dot scales and identify the qualitative distinction between Type I and Type II configurations.
Keywords. $\delta$–shell interaction; singular perturbation; resolvent formula; Kreĭn–type formula; partial–wave decomposition; tunneling splitting
Mathematics Subject Classification (2020). 35J10; 35P15; 47A10; 47F05
1 Introduction
Surface–supported singular perturbations of the Laplacian provide a standard class of solvable Schrödinger operators. A general extension–theoretic framework for solvable models, including point interactions, can be found in Albeverio et al. [2]. For singular interactions supported on hypersurfaces, boundary integral and layer potential approaches were developed by Brasche–Exner–Kuperin–Šeba [7], including an abstract Kreĭn–type resolvent formula in a general measure–valued framework.
For radially symmetric penetrable wall models, Ikebe and Shimada [16] studied the spectral and scattering theory in detail, and Shimada obtained further results on approximation by short–range Hamiltonians [25], low–energy scattering [26], and analytic continuation of the scattering kernel [27]. More recently, essential self–adjointness for systems with several (possibly infinitely many) concentric shells was proved by Albeverio, Kostenko, Malamud, and Neidhardt [3].
The rigorous treatment of $\delta$–interactions supported on a sphere was initiated by Antoine, Gesztesy, and Shabani [4], where the single–shell case was analyzed within an extension–theoretic framework. In the specific setting of finitely many concentric spherical shells with $\delta$–type interactions, Shabani [24] investigated the model by an explicit reduction to radial one–dimensional problems. In particular, for constant shell strengths the matching conditions at the interfaces lead to a finite–dimensional determinant condition in each partial wave, so that the bound state problem can be treated in a fully explicit manner. These works provide closely related precedents for the present study, as they address essentially the same geometric configuration and exhibit, already at the level of partial–wave reduction, the characteristic coupling mechanism induced by multiple shells. While Shabani’s analysis proceeds via a reduction to one–dimensional radial ODEs and finite–dimensional matching conditions in each partial wave (in particular for constant shell strengths), our approach keeps the problem in three dimensions and yields a boundary–integral resolvent representation that remains valid for arbitrary $N$ and nonconstant surface strengths $\alpha_{j}\in L^{\infty}(S_{j};\mathbb{R})$.
Our approach is complementary. Rather than revisiting the operator–theoretic construction, we allow nonconstant strengths $\alpha_{j}\in L^{\infty}(S_{j};\mathbb{R})$ for arbitrary $N$ and derive a concrete boundary integral Kreĭn–type resolvent formula directly in terms of the free Green kernel and single–layer potentials. In the constant two–shell case we then provide a detailed large–separation analysis, including the tuned tunneling splitting on the scale $e^{-\kappa_{0}d}$. This formulation makes explicit the connection between the boundary integral picture and the partial–wave secular equations under rotational symmetry. Related developments on the scattering side for finitely many concentric sphere interactions were later obtained by Hounkonnou, Hounkpe, and Shabani [15].
Building on these works, we consider Schrödinger operators on $\mathbb{R}^{3}$ with finitely many concentric spherical $\delta$–shell interactions. Let $0<R_{1}<R_{2}<\cdots<R_{N}$ and set $S_{j}:=\{x\in\mathbb{R}^{3}:\ |x|=R_{j}\}$ for $j=1,2,\ldots,N$. For each $j$ we allow a bounded measurable surface strength $\alpha_{j}\in L^{\infty}(S_{j};\mathbb{R})$. Formally, the model is
$H_{N}=-\Delta+\sum_{j=1}^{N}\alpha_{j}\,\delta(|x|-R_{j}),$ (1.1)
where $\alpha_{j}$ acts by multiplication on the trace of a function on $S_{j}$. We define $H_{N}$ rigorously by the quadratic form method. At an abstract level, we note that resolvent representations for self–adjoint extensions are well-known. For a wide class of singular perturbations, we can express $(H_{N}-z)^{-1}$ in terms of the free resolvent and an operator–valued boundary term of Kreĭn type. For example, we refer to the extension theoretic framework in [2] and to related formulations for singular interactions in [7]. See also [22] for Kreĭn–type resolvent formulas in a general singular perturbation framework. For $\delta$– and $\delta^{\prime}$–interactions supported on hypersurfaces, including analogues of the Birman–Schwinger principle and a variant of Kreĭn’s formula, we refer to [5]. The main result of this paper is an explicit boundary integral representation for concentric spherical $\delta$–shell interactions, written directly in terms of the free Green kernel and single–layer potentials, without invoking boundary triple machinery.
More precisely, we derive an explicit resolvent formula for $H_{N}$ in terms of layer potentials on the shells. Here $S^{2}=\{\omega\in\mathbb{R}^{3}:\ |\omega|=1\}$ denotes the unit sphere, and this representation yields a boundary operator $K_{N}(z)$ on $\bigoplus_{j=1}^{N}L^{2}(S^{2})$ such that, for $z\in\mathbb{C}\setminus[0,\infty)$, the eigenvalues of $H_{N}$ in $\mathbb{C}\setminus[0,\infty)$ are characterized by the noninvertibility of $K_{N}(z)$.
We then relate this representation to the partial–wave reduction and provide a detailed spectral analysis in the two–shell case. After establishing the NN–shell resolvent framework, we specialize to the two–shell case $N=2$,
$H=-\Delta+\alpha_{1}\,\delta(|x|-R_{1})+\alpha_{2}\,\delta(|x|-R_{2}),\qquad 0<R_{1}<R_{2},$ (1.2)
which already exhibits a nontrivial coupling mechanism between distinct interfaces. For explicit closed formulas and a partial–wave reduction, we further restrict, from that point on, to the rotationally symmetric setting where $\alpha_{1}$ and $\alpha_{2}$ are constants. In this case we obtain a detailed description of the negative spectrum, including a detailed description of the ss–wave eigenvalues $E=-\kappa^{2}<0$. Depending on the signs of $\alpha_{1}$ and $\alpha_{2}$, the ss–wave sector supports zero, one, or two bound states. In the near–decoupling regime of large shell separation $d=R_{2}-R_{1}$, each bound level is close to the corresponding single–shell level. Moreover, when the parameters are tuned so that the two single–shell levels coincide in this limit, say at a common energy $E_{0}=-\kappa_{0}^{2}<0$, where $\kappa_{0}>0$ denotes the decay rate of the corresponding one–shell bound state, the resulting pair of eigenvalues exhibits a tunneling splitting with an exponentially small gap of order $e^{-\kappa_{0}d}$.
A motivation for this analysis comes from semiconductor core–shell nanocrystals. The first observation of size quantization was reported by Ekimov and Onushchenko [12], followed by the effective–mass theory of Efros and Efros [11] and the chemical studies of Rossetti, Nakahara, and Brus [8]. Type I CdSe/ZnS core–shell nanocrystals [14] realize a structure in which both carriers remain confined in the core, while Type II systems with spatially separated electrons and holes were realized later and show distinct excitonic behavior [18, 21]. In the effective–mass picture [11], confinement is produced by band offsets at heterointerfaces, and our $\delta$–shell idealization reflects this mechanism through the signs and positions of the interface parameters.
A related viewpoint appears in the tunneling theory of BenDaniel and Duke [6], where abrupt changes of the effective mass and band edge across a heterojunction are encoded in the BenDaniel–Duke matching conditions. In the effective–mass literature (see, for example, [13]), these conditions are regarded as standard interface rules for very thin heterojunctions. When the intrinsic width of an interface is much smaller than the de Broglie wavelength, its effect can be compressed into a surface term, and $\delta$–type interactions provide a natural effective description. In both the BenDaniel–Duke model and the present $\delta$–shell idealization, the wave function is continuous across the interface, while the normal derivative may jump. In the former case this jump is caused by a change of effective mass, whereas in the latter it is produced by a singular interface potential. In the present work we keep a constant effective mass and absorb interfacial microphysics into effective surface strengths $\alpha_{j}$. Since the bulk potential is taken to be zero away from the shells, a purely repulsive interface ($\alpha>0$) does not by itself produce $L^{2}$ bound states in this idealization, and realistic confinement is primarily due to finite band offsets (finite wells/barriers) and finite interface widths. Accordingly, the discrete eigenvalues analyzed below should be understood as an idealized limit capturing the relevant scales/trends (and tunneling splittings) of effective, possibly quasi–bound, levels rather than as quantitative optical transition energies.
For the rotationally symmetric two–shell setting with constant surface strengths, one can distinguish the Type I and Type II sign patterns familiar from core–shell quantum dots. In this idealized case, Type I systems correspond to $\alpha_{1}<0<\alpha_{2}$ [14], while Type II systems correspond to $\alpha_{1}>0>\alpha_{2}$ [21]. This correspondence is meant only in the constant–strength setting and at the level of sign patterns. Throughout this paper, the Type I/II terminology is used only in the constant case of $\alpha_{j}$.
The paper is organized as follows. In Section 2 we construct $H_{N}$ by quadratic forms and establish basic bounds. In Section 3 we derive the boundary integral resolvent formula and the eigenvalue condition in terms of $K_{N}(z)$. In Sections 4–7 we specialize to two shells with constant couplings, prove that the ground state is an ss–wave, and analyze bound states, level splitting, and tunneling effects for large separation. Appendix A contains a brief order–of–magnitude calibration for representative core–shell quantum dot scales, and Appendix B collects explicit partial–wave matrices used in the rotationally symmetric setting.
2 Form method and operator domain
We fix an integer $N\geq 1$ and radii $0<R_{1}<\cdots<R_{N}$. For each $j\in\{1,\dots,N\}$ we set
$S_{j}:=\{x\in\mathbb{R}^{3}:|x|=R_{j}\}.$
Let $\alpha_{j}$ be a real-valued function on $S_{j}$ such that $\alpha_{j}\in L^{\infty}(S_{j})$. We consider the formal Schrödinger operator
$H_{N}=-\Delta+\sum_{j=1}^{N}\alpha_{j}\,\delta(|x|-R_{j}).$ (2.1)
For $j=1,\dots,N$, let $d\sigma_{j}$ denote the surface measure on $S_{j}:=\{x\in\mathbb{R}^{3}:\ |x|=R_{j}\}$. For $u,v\in H^{1}(\mathbb{R}^{3})$ we define the quadratic form
$h_{N}[u,v]=\int_{\mathbb{R}^{3}}\nabla u\cdot\nabla v\,dx+\sum_{j=1}^{N}\int_{S_{j}}\alpha_{j}(y)\,\bigl(u\upharpoonright_{S_{j}}\bigr)(y)\,\bigl(v\upharpoonright_{S_{j}}\bigr)(y)\,d\sigma_{j}(y).$ (2.2)
Its form domain is $D[h_{N}]=H^{1}(\mathbb{R}^{3})$. Since $\alpha_{j}\in L^{\infty}(S_{j})$ and the trace map $H^{1}(\mathbb{R}^{3})\to L^{2}(S_{j})$ is bounded, each surface term is bounded on $H^{1}(\mathbb{R}^{3})$. Hence $h_{N}$ is closed and lower semibounded and defines a unique self–adjoint operator (which we still denote by $H_{N}$).
To simplify notation, we pull everything back to $S^{2}$. For $u\in H^{1}(\mathbb{R}^{3})$ we define the trace maps $\tau_{j}:H^{1}(\mathbb{R}^{3})\to L^{2}(S^{2})$ by
$(\tau_{j}u)(\omega):=u(R_{j}\omega),\qquad\omega\in S^{2}.$
We also define the pulled–back coefficient
$\widetilde{\alpha}_{j}(\omega):=\alpha_{j}(R_{j}\omega),\qquad\omega\in S^{2}.$
Since $d\sigma_{j}(R_{j}\omega)=R_{j}^{2}\,d\omega$, the surface contribution can be written as
$\int_{S_{j}}\alpha_{j}(y)\,u(y)\,v(y)\,d\sigma_{j}(y)=R_{j}^{2}\int_{S^{2}}\widetilde{\alpha}_{j}(\omega)\,(\tau_{j}u)(\omega)\,(\tau_{j}v)(\omega)\,d\omega.$
We set
$\Omega_{1}$ $=$ $\bigl\{|x|<R_{1}\bigr\},$
$\Omega_{k}$ $=$ $\bigl\{R_{k-1}<|x|<R_{k}\bigr\},\qquad k=2,\dots,N,$
$\Omega_{N+1}$ $=$ $\bigl\{|x|>R_{N}\bigr\}.$
Moreover, the operator domain $D(H_{N})$ is characterized by the usual interface rules: $u$ is continuous across each $S_{j}$, and the radial derivative satisfies
$\partial_{r}u(R_{j}+0,\omega)-\partial_{r}u(R_{j}-0,\omega)=\widetilde{\alpha}_{j}(\omega)\,u(R_{j},\omega),\qquad\omega\in S^{2},\quad j=1,\dots,N.$ (2.3)
More precisely,
$D(H_{N})$ $=$ $\Bigl\{u\in L^{2}(\mathbb{R}^{3})\ \Bigm|\ u\upharpoonright_{\Omega_{k}}\in H^{2}(\Omega_{k})\ \text{for }k=1,\dots,N+1,$ (2.4)
[…]
On each region $\Omega_{k}$ the operator acts as $(H_{N}u)\upharpoonright_{\Omega_{k}}=-\Delta(u|_{\Omega_{k}})$.
We begin with a simple observation excluding positive eigenvalues.
Theorem 1.
Let $H_{N}$ be the Schrödinger operator on $\mathbb{R}^{3}$ with finitely many concentric spherical $\delta$–shell interactions (2.1), where $\alpha_{j}\in L^{\infty}(S_{j};\mathbb{R})$. Then $H_{N}$ has no eigenvalues in $(0,\infty)$.
Proof.
Fix $E=k^{2}>0$ and assume that $H_{N}u=Eu$ for some $u\in L^{2}(\mathbb{R}^{3})$. Then $u\in D(H_{N})$. In particular, $u\in H^{2}(\Omega_{k})$ on each region $\Omega_{k}$ away from the shells, and the traces $u|_{S_{j}}$ and $\partial_{r}^{\pm}u\upharpoonright_{S_{j}}$ are well-defined.
For almost every $r>0$, the function $u(r,\cdot)$ belongs to $L^{2}(S^{2})$. We therefore expand
$u(r,\omega)=\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}\frac{f_{\ell m}(r)}{r}\,Y_{\ell m}(\omega),\qquad f_{\ell m}(r)=r\int_{S^{2}}u(r,\omega)\overline{Y_{\ell m}(\omega)}\,d\omega,$ (2.5)
where $\{Y_{\ell m}\}$ is an orthonormal basis of $L^{2}(S^{2})$. See [1]. By Parseval’s identity,
$r^{2}\int_{S^{2}}|u(r,\omega)|^{2}\,d\omega=\sum_{\ell,m}|f_{\ell m}(r)|^{2}\quad\text{for a.e.\ }r>0,$ (2.6)
and consequently
$\int_{|x|>R_{N}}|u(x)|^{2}\,dx=\int_{R_{N}}^{\infty}\sum_{\ell,m}|f_{\ell m}(r)|^{2}\,dr.$ (2.7)
Away from the shells $r=R_{1},\dots,R_{N}$ the equation reduces to $(-\Delta-k^{2})u=0$. Substituting (2.5) (equivalently, projecting $(-\Delta-k^{2})u=0$ onto the spherical harmonics) shows that each $f_{\ell m}$ satisfies
$f_{\ell m}^{\prime\prime}(r)+\frac{\ell(\ell+1)}{r^{2}}f_{\ell m}(r)=k^{2}f_{\ell m}(r)$ (2.8)
on every interval not containing any $R_{j}$.
For $r>R_{N}$ the general solution of (2.8) can be written as
$f_{\ell m}(r)=A_{\ell m}\,r\,h^{(1)}_{\ell}(kr)+B_{\ell m}\,r\,h^{(2)}_{\ell}(kr),$ (2.9)
where $h^{(1)}_{\ell}$ and $h^{(2)}_{\ell}$ are the spherical Hankel functions. Using the asymptotics
$r\,h^{(1)}_{\ell}(kr)=\frac{e^{ikr}}{ik}+O(r^{-1}),\qquad r\,h^{(2)}_{\ell}(kr)=-\frac{e^{-ikr}}{ik}+O(r^{-1}),\quad r\to\infty,$
we obtain
$f_{\ell m}(r)=p_{\ell m}e^{ikr}+q_{\ell m}e^{-ikr}+O(r^{-1}),\qquad r\to\infty,$ (2.10)
for suitable constants $p_{\ell m},q_{\ell m}$.
Assume that $u$ does not vanish identically in the exterior region $(R_{N},\infty)$. Then there exists $(\ell_{0},m_{0})$ such that $f_{\ell_{0}m_{0}}$ is not identically zero on $(R_{N},\infty)$. Write $f=f_{\ell_{0}m_{0}}$ and $g(r)=pe^{ikr}+qe^{-ikr}$ with $M:=|p|^{2}+|q|^{2}>0$. By (2.10) there exist $C>0$ and $R_{0}>R_{N}$ such that $|f(r)-g(r)|\leq C/r$ for $r\geq R_{0}$.
For every $R\geq R_{0}$ we have
$\int_{R}^{R+\pi/k}|g(r)|^{2}\,dr=\int_{R}^{R+\pi/k}\Bigl(M+2\text{{\rm Re}}(p\overline{q}e^{2ikr})\Bigr)\,dr=\frac{\pi}{k}M,$
since $\int_{R}^{R+\pi/k}e^{2ikr}\,dr=0$. Moreover, using $|a+b|^{2}\geq\frac{1}{2}|a|^{2}-|b|^{2}$, we obtain
$\int_{R}^{R+\pi/k}|f(r)|^{2}\,dr$ $\geq$ $\frac{1}{2}\int_{R}^{R+\pi/k}|g(r)|^{2}\,dr-\int_{R}^{R+\pi/k}|f(r)-g(r)|^{2}\,dr$
$\geq$ $\frac{\pi}{2k}M-\frac{C^{2}\pi}{kR^{2}}.$
Choosing $R$ so large that $C^{2}/R^{2}\leq M/4$, we obtain
$\int_{R}^{R+\pi/k}|f(r)|^{2}\,dr\geq\frac{\pi}{4k}M=:c>0.$
Summing over the disjoint intervals $R+n\pi/k,R+(n+1)\pi/k]$ yields
$\int_{R}^{\infty}|f(r)|^{2}\,dr=\infty.$
Using (2.7) and $\sum_{\ell,m}|f_{\ell m}(r)|^{2}\geq|f(r)|^{2}$, we conclude that
$\int_{|x|>R_{N}}|u(x)|^{2}\,dx=\infty,$
contradicting $u\in L^{2}(\mathbb{R}^{3})$. Therefore $u\equiv 0$ for $r>R_{N}$.
By continuity, $u|_{S_{N}}=0$, and since $u$ vanishes identically in the exterior region we also have $\partial_{r}^{+}u|_{S_{N}}=0$. The jump condition (2.3) on $S_{N}$ implies $\partial_{r}^{-}u|_{S_{N}}=0$. Hence $f_{\ell m}(R_{N})=0$ for all $\ell,m$. Moreover, since $u\in H^{2}(\Omega_{N})\cap H^{2}(\Omega_{N+1})$, we have
$f^{\prime}_{\ell m}(R_{N}\pm 0)=R_{N}\int_{S^{2}}\partial_{r}^{\pm}u(R_{N},\omega)\overline{Y_{\ell m}(\omega)}\,d\omega,$
and thus $f^{\prime}_{\ell m}(R_{N})=0$ for all $\ell,m$. Uniqueness for the second–order ODE (2.8) yields $f_{\ell m}\equiv 0$ on $(R_{N-1},R_{N})$. Repeating the argument across the remaining shells shows that $u\equiv 0$ on $\mathbb{R}^{3}$, a contradiction. Thus there are no positive eigenvalues. ∎
When the surface strengths are constants, rotational symmetry allows a partial–wave decomposition. At the zero–energy threshold $E=0$, each angular momentum channel reduces to a finite–dimensional algebraic condition on the shell values, which can be written explicitly.
Theorem 2 (Zero–energy threshold).
Fix an integer $N\geq 1$ and radii $0<R_{1}<\cdots<R_{N}$. Assume that the surface strengths are constants $\alpha_{1},\dots,\alpha_{N}\in\mathbb{R}$. Let $\ell\in\{0,1,2,\dots\}$ be the angular–momentum index.
(i) In the ss–wave case $\ell=0$, the point $E=0$ is not an $L^{2}$ eigenvalue for any finite $\alpha_{1},\dots,\alpha_{N}$.
(ii) For each fixed $\ell\geq 1$, define the $N\times N$ matrix $A_{\ell}=(a^{(\ell)}_{ij})_{i,j=1}^{N}$ by
$a^{(\ell)}_{ij}:=\delta_{ij}+\frac{\alpha_{j}}{2\ell+1}\,\frac{R_{\min(i,j)}^{\ell+1}}{R_{\max(i,j)}^{\ell}},\qquad i,j=1,\dots,N,$ (2.11)
where $R_{\min(i,j)}:=\min(R_{i},R_{j})$ and $R_{\max(i,j)}:=\max(R_{i},R_{j})$. Then $E=0$ is an $L^{2}$ eigenvalue in the $\ell$–th partial wave if and only if
$\det A_{\ell}=0.$ (2.12)
Moreover, the multiplicity of $E=0$ contributed by the $\ell$–th partial wave equals $(2\ell+1)\dim\text{\rm Ker}A_{\ell}$.
Proof.
Fix $\ell\geq 0$ and write
$\psi(x)=\frac{u(r)}{r}Y_{\ell m}(\omega),\qquad r=|x|,\ \omega=x/r,$
so that on each region the radial function satisfies
$u^{\prime\prime}(r)+\frac{\ell(\ell+1)}{r^{2}}u(r)=E\,u(r),\qquad r\neq R_{1},\dots,R_{N}.$
Moreover, since $\{Y_{\ell m}\}$ is an orthonormal basis of $L^{2}(S^{2})$, we have $\|\psi\|_{L^{2}(\mathbb{R}^{3})}^{2}=\int_{0}^{\infty}|u(r)|^{2}\,dr$. The $\delta$–shell interface conditions for $\psi$ are equivalent to the continuity of $u$ at each $R_{j}$ and to the jump conditions
$u^{\prime}(R_{j}+0)-u^{\prime}(R_{j}-0)=\alpha_{j}\,u(R_{j}),\qquad j=1,\dots,N.$
We consider $E=0$.
(i) Let $\ell=0$. Then $u^{\prime\prime}(r)=0$ on each interval, so on the exterior region $r>R_{N}$ we have $u(r)=ar+b$. Since $\int_{R_{N}}^{\infty}|u(r)|^{2}\,dr<\infty$, it follows that $a=b=0$ and hence $u\equiv 0$ on $(R_{N},\infty)$. By continuity, $u(R_{N})=0$, and the jump condition at $R_{N}$ yields $u^{\prime}(R_{N}-0)=u^{\prime}(R_{N}+0)=0$. Since $u^{\prime\prime}(r)=0$ on $(R_{N-1},R_{N})$, these two boundary conditions imply that $u\equiv 0$ on $(R_{N-1},R_{N})$. Iterating this argument across the remaining shells, we obtain $u\equiv 0$ on $(0,\infty)$, hence $\psi\equiv 0$, a contradiction. Therefore $E=0$ cannot be an $L^{2}$ eigenvalue in the ss–wave sector.
(ii) Assume $\ell\geq 1$. Consider the homogeneous equation
$w^{\prime\prime}(r)+\frac{\ell(\ell+1)}{r^{2}}w(r)=0\qquad(r\neq s),$
whose independent solutions are $r^{\ell+1}$ and $r^{-\ell}$. Let $r_{<}:=\min(r,s)$ and $r_{>}:=\max(r,s)$ and define
$G_{\ell}(r,s):=\frac{1}{2\ell+1}\,r_{<}^{\ell+1}r_{>}^{-\ell}.$ (2.13)
Then $G_{\ell}(\cdot,s)$ is continuous, solves the homogeneous equation for $r\neq s$, decays like $r^{-\ell}$ as $r\to\infty$, is regular at $r=0$, and satisfies $G^{\prime}_{\ell}(s+0,s)-G^{\prime}_{\ell}(s-0,s)=-1$. Consequently, for any constants $c_{1},\dots,c_{N}$, the function
$w(r):=-\sum_{j=1}^{N}c_{j}\,G_{\ell}(r,R_{j})$
is continuous on $(0,\infty)$, solves the homogeneous equation away from the shells, and has derivative jumps $w^{\prime}(R_{j}+0)-w^{\prime}(R_{j}-0)=c_{j}$.
Let $\psi$ be an $L^{2}$ solution at $E=0$ in the $\ell$–th partial wave and set $U_{j}:=u(R_{j})$. Define
$\widetilde{u}(r):=-\sum_{j=1}^{N}\alpha_{j}U_{j}\,G_{\ell}(r,R_{j}).$
By the preceding properties of $G_{\ell}$, the function $\widetilde{u}$ is continuous, solves the radial equation for $r\neq R_{j}$, and satisfies $\widetilde{u}^{\prime}(R_{j}+0)-\widetilde{u}^{\prime}(R_{j}-0)=\alpha_{j}U_{j}$. Hence $v:=u-\widetilde{u}$ solves the homogeneous equation on $(0,\infty)$, is continuous with no derivative jumps at any $R_{j}$, is regular at 0, and belongs to $L^{2}(0,\infty)$. Therefore $v\equiv 0$, and thus $u=\widetilde{u}$. Evaluating at $r=R_{i}$ gives, for $i=1,\dots,N$,
$U_{i}=-\sum_{j=1}^{N}\alpha_{j}U_{j}\,G_{\ell}(R_{i},R_{j})=-\sum_{j=1}^{N}\alpha_{j}U_{j}\,\frac{1}{2\ell+1}\,\frac{R_{\min(i,j)}^{\ell+1}}{R_{\max(i,j)}^{\ell}}.$
This is exactly the linear system $A_{\ell}U=0$ with $A_{\ell}$ as in (2.11). Hence a nontrivial $L^{2}$ solution exists if and only if $\text{\rm Ker}A_{\ell}\neq\{0\}$, equivalently (2.12) holds.
Finally, the radial equation and the interface conditions do not depend on $m$, so each independent radial solution produces $(2\ell+1)$ linearly independent eigenfunctions by varying $m$. This gives the multiplicity formula $(2\ell+1)\dim\text{\rm Ker}A_{\ell}$. ∎
Remark 1.
For $N=2$ with $0<R_{1}<R_{2}$ and constants $\alpha_{1},\alpha_{2}\in\mathbb{R}$, the matrix $A_{\ell}$ in (2.11) takes the form
$A_{\ell}=\left(\begin{array}[]{cc}1+\dfrac{\alpha_{1}R_{1}}{2\ell+1}&\dfrac{\alpha_{2}}{2\ell+1}\dfrac{R_{1}^{\ell+1}}{R_{2}^{\ell}}\\[8.53581pt] \dfrac{\alpha_{1}}{2\ell+1}\dfrac{R_{1}^{\ell+1}}{R_{2}^{\ell}}&1+\dfrac{\alpha_{2}R_{2}}{2\ell+1}\end{array}\right).$
Thus the condition $\det A_{\ell}=0$ becomes
$\alpha_{1}\alpha_{2}\,R_{1}^{2\ell+2}R_{2}^{-2\ell}=\bigl(\alpha_{1}R_{1}+2\ell+1\bigr)\bigl(\alpha_{2}R_{2}+2\ell+1\bigr).$
3 Resolvent Formula
In this section we derive an explicit resolvent formula for the NN–shell Hamiltonian by using the free Green kernel and single–layer potentials. We do not use boundary triples or more abstract extension theory. Throughout, we take the principal branch so that $\text{{\rm Im}}\sqrt{z}>0$, and we set $k=\sqrt{z}$.
Fix $z\in\mathbb{C}\setminus[0,\infty)$ and write
$G_{z}(x,y)=\frac{e^{ik|x-y|}}{4\pi|x-y|},\qquad(R_{0}(z)f)(x)=\int_{\mathbb{R}^{3}}G_{z}(x,y)f(y)\,dy.$
For $u\in H^{1}(\mathbb{R}^{3})$ we define the pulled–back trace maps $\tau_{j}:H^{1}(\mathbb{R}^{3})\to L^{2}(S^{2})$ by
$(\tau_{j}u)(\omega)=u(R_{j}\omega),\qquad\omega\in S^{2}.$
Under the parametrization $y=R_{j}\omega$, $\omega\in S^{2}$, one has $d\sigma_{j}(y)=R_{j}^{2}\,d\omega$, hence
$\int_{S_{j}}|u(y)|^{2}\,d\sigma_{j}(y)=R_{j}^{2}\int_{S^{2}}|(\tau_{j}u)(\omega)|^{2}\,d\omega.$
We also write
$\widetilde{\alpha}_{j}(\omega)=\alpha_{j}(R_{j}\omega),\qquad\omega\in S^{2}.$
For $\varphi\in L^{2}(S^{2})$ we define the single–layer potentials supported on $S_{j}$ by
$(\Gamma_{j}(z)\varphi)(x)=\int_{S^{2}}G_{z}\bigl(x,R_{j}\omega\bigr)\,\varphi(\omega)\,d\omega,\qquad x\in\mathbb{R}^{3}.$ (3.1)
We introduce the boundary Hilbert space
$\mathcal{K}_{N}:=\bigoplus_{j=1}^{N}L^{2}(S^{2}),$ (3.2)
[…]
$m_{ij}(z):L^{2}(S^{2})\to L^{2}(S^{2}),\qquad(m_{ij}(z)\varphi)(\omega)=\int_{S^{2}}G_{z}\bigl(R_{i}\omega,R_{j}\omega^{\prime}\bigr)\,\varphi(\omega^{\prime})\,d\omega^{\prime}.$ (3.3)
We also write $\Gamma(z)=(\Gamma_{1}(z),\dots,\Gamma_{N}(z))$ as a column operator
$\Gamma(z):\mathcal{K}_{N}\to L^{2}(\mathbb{R}^{3}),\qquad\Gamma(z)\,{}^{t}(\varphi_{1},\dots,\varphi_{N})=\sum_{j=1}^{N}\Gamma_{j}(z)\varphi_{j}.$
Finally we define a bounded diagonal operator $\Theta$ on $\mathcal{K}_{N}$ by
$(\Theta\,{}^{t}(\varphi_{1},\dots,\varphi_{N}))_{j}(\omega)=R_{j}^{2}\,\widetilde{\alpha}_{j}(\omega)\,\varphi_{j}(\omega),\qquad j=1,\dots,N,$ (3.4)
and set
$K_{N}(z)=I+m(z)\Theta.$ (3.5)
Lemma 3.
Let $R>0$ and $z\in\mathbb{C}\setminus[0,\infty)$ with $\text{{\rm Im}}\sqrt{z}>0$. Define
$(\widetilde{\Gamma}(z)\varphi)(x)=\int_{S^{2}}G_{z}(x,R\omega)\,\varphi(\omega)\,d\omega,\qquad\varphi\in L^{2}(S^{2}).$
Then $\widetilde{\Gamma}(z):L^{2}(S^{2})\to L^{2}(\mathbb{R}^{3})$ is a Hilbert–Schmidt operator.
Proof.
The Hilbert–Schmidt norm is
$\|\widetilde{\Gamma}(z)\|_{\mathrm{HS}}^{2}=\int_{\mathbb{R}^{3}}\int_{S^{2}}\bigl|G_{z}(x,R\omega)\bigr|^{2}\,d\omega\,dx.$
Let $c=\text{{\rm Im}}\sqrt{z}>0$. Since
$\bigl|G_{z}(x,y)\bigr|\leq\frac{1}{4\pi}\,\frac{e^{-c|x-y|}}{|x-y|}\qquad(x\neq y),$
we obtain
$\|\widetilde{\Gamma}(z)\|_{\mathrm{HS}}^{2}\leq\frac{1}{(4\pi)^{2}}\int_{\mathbb{R}^{3}}\int_{S^{2}}\frac{e^{-2c|x-R\omega|}}{|x-R\omega|^{2}}\,d\omega\,dx.$
Write $r=|x|$ and let $\theta$ be the angle between $x/|x|$ and $\omega$. Then $\rho=|x-R\omega|=\sqrt{r^{2}+R^{2}-2rR\cos\theta}$, and $d\omega=2\pi\sin\theta\,d\theta$ with $\sin\theta\,d\theta=(\rho/(rR))\,d\rho$. Hence
$\int_{S^{2}}\frac{e^{-2c|x-R\omega|}}{|x-R\omega|^{2}}\,d\omega=\frac{2\pi}{rR}\int_{|r-R|}^{r+R}\frac{e^{-2c\rho}}{\rho}\,d\rho.$
Using $dx=4\pi r^{2}\,dr$ and Fubini–Tonelli, we get
$\|\widetilde{\Gamma}(z)\|_{\mathrm{HS}}^{2}$ $\leq$ $\frac{1}{(4\pi)^{2}}\int_{0}^{\infty}\left(\frac{2\pi}{rR}\int_{|r-R|}^{r+R}\frac{e^{-2c\rho}}{\rho}\,d\rho\right)4\pi r^{2}\,dr$
$=$ $\frac{1}{2R}\int_{0}^{\infty}r\int_{|r-R|}^{r+R}\frac{e^{-2c\rho}}{\rho}\,d\rho\,dr.$
The domain is
$D=\Bigl\{(r,\rho):\ r\geq 0,\ |r-R|\leq\rho\leq r+R\Bigr\}=\Bigl\{(r,\rho):\ \rho\geq 0,\ |R-\rho|\leq r\leq R+\rho\Bigr\}.$
Exchanging the order of integration yields
$\|\widetilde{\Gamma}(z)\|_{\mathrm{HS}}^{2}$ $\leq$ $\frac{1}{2R}\int_{0}^{\infty}\frac{e^{-2c\rho}}{\rho}\left(\int_{|R-\rho|}^{R+\rho}r\,dr\right)d\rho=\int_{0}^{\infty}e^{-2c\rho}\,d\rho=\frac{1}{2c}<\infty.$
Lemma 4.
Let $z\in\mathbb{C}\setminus[0,\infty)$ and $\text{{\rm Im}}\sqrt{z}>0$. Then the following statements hold.
(a) Each $\Gamma_{j}(z):L^{2}(S^{2})\to L^{2}(\mathbb{R}^{3})$ is a Hilbert–Schmidt operator, and in particular $\Gamma(z)$ is Hilbert–Schmidt.
(b) For every $i,j$ the operator $m_{ij}(z)$ is bounded on $L^{2}(S^{2})$ and one has $m_{ij}(\bar{z})=m_{ji}(z)^{\ast}$. Moreover, the identity $\tau_{i}\Gamma_{j}(z)=m_{ij}(z)$ holds in $L^{2}(S^{2})$.
(c) The adjoint satisfies $\Gamma_{j}(\bar{z})^{\ast}=\tau_{j}R_{0}(z)$ as an operator from $L^{2}(\mathbb{R}^{3})$ to $L^{2}(S^{2})$.
Proof.
(a) follows from Lemma 3 with $R=R_{j}$.
(b) The operator $m_{ij}(z)$ is an integral operator with kernel $K_{ij}^{z}(\omega,\omega^{\prime})=G_{z}(R_{i}\omega,R_{j}\omega^{\prime})$. Let $c=\text{{\rm Im}}\sqrt{z}>0$. Then
$\bigl|K_{ij}^{z}(\omega,\omega^{\prime})\bigr|\leq\frac{1}{4\pi}\,\frac{e^{-c|R_{i}\omega-R_{j}\omega^{\prime}|}}{|R_{i}\omega-R_{j}\omega^{\prime}|}\qquad(\omega\neq\omega^{\prime}).$
For fixed $\omega$, the map $\omega^{\prime}\mapsto|R_{i}\omega-R_{j}\omega^{\prime}|^{-1}$ is integrable on $S^{2}$, and similarly with $\omega$ and $\omega^{\prime}$ interchanged. Hence the Schur test yields boundedness of $m_{ij}(z)$ on $L^{2}(S^{2})$. The symmetry $G_{\bar{z}}(x,y)=\overline{G_{z}(y,x)}$ implies $m_{ij}(\bar{z})=m_{ji}(z)^{\ast}$.
To prove $\tau_{i}\Gamma_{j}(z)=m_{ij}(z)$, let $\phi\in L^{2}(S^{2})$. The estimate above implies absolute integrability on $S^{2}\times S^{2}$, hence Fubini–Tonelli applies and, for a.e. $\omega\in S^{2}$,
$(\tau_{i}\Gamma_{j}(z)\phi)(\omega)=\int_{S^{2}}G_{z}(R_{i}\omega,R_{j}\omega^{\prime})\,\phi(\omega^{\prime})\,d\omega^{\prime}=(m_{ij}(z)\phi)(\omega).$
(c) Let $f\in L^{2}(\mathbb{R}^{3})$ and $\phi\in L^{2}(S^{2})$. Using Fubini–Tonelli (justified by Cauchy–Schwarz and (a)), we compute
$(\Gamma_{j}(\bar{z})\phi,f)_{L^{2}(\mathbb{R}^{3})}$ $=$ $\int_{\mathbb{R}^{3}}\int_{S^{2}}G_{\bar{z}}(x,R_{j}\omega)\,\phi(\omega)\,\overline{f(x)}\,d\omega\,dx$
$=$ $\int_{S^{2}}\phi(\omega)\,\overline{\int_{\mathbb{R}^{3}}\overline{G_{\bar{z}}(x,R_{j}\omega)}\,f(x)\,dx}\,d\omega$
$=$ $\int_{S^{2}}\phi(\omega)\,\overline{\int_{\mathbb{R}^{3}}G_{z}(R_{j}\omega,x)\,f(x)\,dx}\,d\omega=(\phi,\ \tau_{j}R_{0}(z)f)_{L^{2}(S^{2})}.$
Hence $\Gamma_{j}(\bar{z})^{\ast}=\tau_{j}R_{0}(z)$. ∎
We recall the quadratic form $h_{N}$ from Section 2,
$h_{N}[u,v]=\int_{\mathbb{R}^{3}}\nabla u\cdot\overline{\nabla v}\,dx+\sum_{j=1}^{N}\int_{S_{j}}\alpha_{j}(x)\,u(x)\,\overline{v(x)}\,d\sigma_{j}(x),\qquad D[h_{N}]=H^{1}(\mathbb{R}^{3}),$ (3.6)
which defines a self–adjoint operator $H_{N}$ in $L^{2}(\mathbb{R}^{3})$.
While Kreĭn–type resolvent formulas are available in abstract extension theory, we present here a self–contained boundary integral derivation tailored to concentric spheres, which yields an explicit $N\times N$ operator matrix $K_{N}(z)$ on $\mathcal{K}_{N}=\bigoplus_{j=1}^{N}L^{2}(S^{2})$ and is directly compatible with the subsequent partial–wave reduction.
We are now in a position to state the resolvent formula for $H_{N}$, which is one of the main results of this paper.
Theorem 5.
Let $H_{N}$ be as above and let $z\in\mathbb{C}\setminus[0,\infty)$ with $\text{{\rm Im}}\sqrt{z}>0$. Assume that $K_{N}(z)=I+m(z)\Theta$ on $\bigoplus_{j=1}^{N}L^{2}(S^{2})$ is invertible, where $\Theta$ is given in (3.4). Define a bounded operator $R(z)$ on $L^{2}(\mathbb{R}^{3})$ by
$R(z)=R_{0}(z)-\Gamma(z)\,\Theta\,K_{N}(z)^{-1}\Gamma(\bar{z})^{\ast}.$ (3.7)
Then $R(z)=(H_{N}-z)^{-1}$, in particular $z\in\rho(H_{N})$, and the resolvent difference $(H_{N}-z)^{-1}-R_{0}(z)$ is a trace class operator.
Proof.
Assume that $K_{N}(z)$ is invertible and define $R(z)$ by (3.7). Let $f\in L^{2}(\mathbb{R}^{3})$ and set
$u:=R(z)f.$ (3.8)
We show that $u\in D(H_{N})$ and $(H_{N}-z)u=f$.
We first verify that $u\in H^{1}(\mathbb{R}^{3})$. Since $R_{0}(z):L^{2}(\mathbb{R}^{3})\to H^{2}(\mathbb{R}^{3})\subset H^{1}(\mathbb{R}^{3})$, we have $R_{0}(z)f\in H^{1}(\mathbb{R}^{3})$. Each trace map $\tau_{j}:H^{1}(\mathbb{R}^{3})\to H^{1/2}(S^{2})$ is bounded, and we use the continuous embedding $H^{1/2}(S^{2})\hookrightarrow L^{2}(S^{2})$ to regard $\tau_{j}$ as a bounded operator into $L^{2}(S^{2})$. Hence $\tau_{j}:H^{1}(\mathbb{R}^{3})\to L^{2}(S^{2})$ is bounded, and its adjoint $\tau_{j}^{\ast}:L^{2}(S^{2})\to H^{-1}(\mathbb{R}^{3})$ is bounded. For $z\in\mathbb{C}\setminus[0,\infty)$ the free resolvent extends boundedly as $R_{0}(z):H^{-1}(\mathbb{R}^{3})\to H^{1}(\mathbb{R}^{3})$. By Lemma 4, one has the identity $R_{0}(z)\tau_{j}^{\ast}=\Gamma_{j}(z)$. Since $R_{0}(z):H^{-1}(\mathbb{R}^{3})\to H^{1}(\mathbb{R}^{3})$ is bounded, it follows that $\Gamma_{j}(z)$ maps $L^{2}(S^{2})$ boundedly into $H^{1}(\mathbb{R}^{3})$. Hence (3.8) implies $u\in H^{1}(\mathbb{R}^{3})$. Set
[…]
Using again Lemma 4, namely $\Gamma_{j}(\bar{z})^{\ast}=\tau_{j}R_{0}(z)$, we obtain
$u=R_{0}(z)f-\sum_{j=1}^{N}\Gamma_{j}(z)\,(\Theta g)_{j},\qquad(\Theta g)_{j}(\omega)=R_{j}^{2}\,\widetilde{\alpha}_{j}(\omega)\,g_{j}(\omega).$
Taking traces and using $\tau_{i}\Gamma_{j}(z)=m_{ij}(z)$ from Lemma 4(b), we obtain
$\tau_{i}u=\tau_{i}R_{0}(z)f-\sum_{j=1}^{N}m_{ij}(z)\,(\Theta g)_{j},\qquad i=1,\dots,N.$
[…]
$\tau u=\Gamma(\bar{z})^{\ast}f-m(z)\Theta g.$
Since $K_{N}(z)=I+m(z)\Theta$ and $K_{N}(z)g=\Gamma(\bar{z})^{\ast}f$, we obtain
$\tau u=K_{N}(z)g-m(z)\Theta g=g.$
Let $\varphi\in H^{1}(\mathbb{R}^{3})$ and let $H_{0}:=-\Delta$ be the free Schrödinger operator. Since $(H_{0}-z)R_{0}(z)f=f$, we have
$(\nabla R_{0}(z)f,\nabla\varphi)_{L^{2}(\mathbb{R}^{3})}-z(R_{0}(z)f,\varphi)_{L^{2}(\mathbb{R}^{3})}=(f,\varphi)_{L^{2}(\mathbb{R}^{3})}.$
Moreover, for $h\in L^{2}(S^{2})$ one has $(H_{0}-z)\Gamma_{j}(z)h=\tau_{j}^{\ast}h$ in $H^{-1}(\mathbb{R}^{3})$, hence
$(\nabla\Gamma_{j}(z)h,\nabla\varphi)_{L^{2}(\mathbb{R}^{3})}-z(\Gamma_{j}(z)h,\varphi)_{L^{2}(\mathbb{R}^{3})}=(h,\tau_{j}\varphi)_{L^{2}(S^{2})}.$
Applying this with $h=(\Theta g)_{j}$ and summing over $j$, we obtain
$(\nabla u,\nabla\varphi)_{L^{2}(\mathbb{R}^{3})}-z(u,\varphi)_{L^{2}(\mathbb{R}^{3})}=(f,\varphi)_{L^{2}(\mathbb{R}^{3})}-\sum_{j=1}^{N}((\Theta g)_{j},\tau_{j}\varphi)_{L^{2}(S^{2})}.$
On the other hand, by (3.6) and $d\sigma_{j}=R_{j}^{2}\,d\omega$,
$h_{N}[u,\varphi]-z(u,\varphi)_{L^{2}(\mathbb{R}^{3})}=(\nabla u,\nabla\varphi)_{L^{2}(\mathbb{R}^{3})}-z(u,\varphi)_{L^{2}(\mathbb{R}^{3})}+\sum_{j=1}^{N}((\Theta\,\tau u)_{j},\tau_{j}\varphi)_{L^{2}(S^{2})}.$
Since $\tau u=g$, we have $\Theta\,\tau u=\Theta g$, and hence
$h_{N}[u,\varphi]-z(u,\varphi)_{L^{2}(\mathbb{R}^{3})}=(f,\varphi)_{L^{2}(\mathbb{R}^{3})}\qquad(\forall\,\varphi\in H^{1}(\mathbb{R}^{3})).$
Since $h_{N}$ is closed and lower semibounded, by the representation theorem (with $w:=f+zu\in L^{2}(\mathbb{R}^{3})$) this implies that $u\in D(H_{N})$ and $(H_{N}-z)u=f$. Therefore $R(z)=(H_{N}-z)^{-1}$.
Finally, by Lemma 4(a), both $\Gamma(z)$ and $\Gamma(\bar{z})^{\ast}$ are Hilbert–Schmidt operators. Since $\Theta$ and $K_{N}(z)^{-1}$ are bounded, the composition $\Gamma(z)\Theta K_{N}(z)^{-1}\Gamma(\bar{z})^{\ast}$ is a trace class operator, by the ideal property of the Schatten–von Neumann classes. Consequently, the resolvent difference $(H_{N}-z)^{-1}-R_{0}(z)$ is trace class. ∎
Lemma 6.
For each $z\in\mathbb{C}\setminus[0,\infty)$ the operator $m(z)$ acts boundedly on $\bigoplus_{j=1}^{N}L^{2}(S^{2})$, depends analytically on $z$, and is compact. Hence $K_{N}(z)=I+m(z)\Theta$ is an analytic Fredholm family of index zero. In particular, if $K_{N}(z_{0})$ is invertible for some $z_{0}\in\mathbb{C}\setminus[0,\infty)$, then $K_{N}(z)^{-1}$ exists and depends meromorphically on $z$, and the set of points where $K_{N}(z)$ is not invertible is discrete. These points coincide with the poles of the resolvent of $H_{N}$; see Theorem 5 together with the resolvent identity (3.7).
Proof.
Each block $m_{ij}(z)$ is an integral operator with kernel $G_{z}(R_{i}\omega,R_{j}\omega^{\prime})$ on the compact set $S^{2}\times S^{2}$.
If $i\neq j$, then $|R_{i}\omega-R_{j}\omega^{\prime}|\geq|R_{i}-R_{j}|>0$, so the kernel is continuous on $S^{2}\times S^{2}$ for $\text{{\rm Im}}\sqrt{z}>0$. Hence $m_{ij}(z)$ is Hilbert–Schmidt and in particular compact.
For $i=j$, we write
$m_{ii}(z)=m_{ii}(0)+\bigl(m_{ii}(z)-m_{ii}(0)\bigr).$
The difference $G_{z}-G_{0}$ has a removable singularity at the diagonal, since
$\frac{e^{i\sqrt{z}r}-1}{4\pi r}=\frac{i\sqrt{z}}{4\pi}+O(r)\qquad(r\to 0).$
Thus the kernel of $m_{ii}(z)-m_{ii}(0)$ extends continuously to $S^{2}\times S^{2}$, so $m_{ii}(z)-m_{ii}(0)$ is Hilbert–Schmidt and hence compact. On the other hand, $m_{ii}(0)$ is the Laplace single–layer operator on $S^{2}$. It maps $L^{2}(S^{2})$ continuously into $H^{1}(S^{2})$, and the embedding $H^{1}(S^{2})\hookrightarrow L^{2}(S^{2})$ is compact. Therefore $m_{ii}(0)$ is compact on $L^{2}(S^{2})$, and so is $m_{ii}(z)$.
Hence all blocks $m_{ij}(z)$ are compact and $m(z)$ is compact on $\mathcal{K}_{N}$. Analyticity in $z$ follows from the explicit kernel dependence for $\text{{\rm Im}}\sqrt{z}>0$. Since $\Theta$ is bounded, $m(z)\Theta$ is compact and $K_{N}(z)=I+m(z)\Theta$ is an analytic Fredholm family of index zero. The remaining statements follow from the analytic Fredholm theorem (applied once $K_{N}(z_{0})$ is invertible for some $z_{0}\in\mathbb{C}\setminus[0,\infty)$) together with Theorem 5 and (3.7). ∎
The following proposition characterizes the eigenvalues of $H_{N}$ in terms of the boundary operator $K_{N}(z)$.
Proposition 7.
Let $z\in\mathbb{C}\setminus[0,\infty)$. Then $z$ belongs to the point spectrum of $H_{N}$ if and only if $K_{N}(z)$ is not invertible. Moreover, if $K_{N}(z)$ is not invertible, then
$\dim\text{\rm Ker}(H_{N}-z)=\dim\text{\rm Ker}K_{N}(z).$
Proof.
If $K_{N}(z)$ is invertible, then Theorem 5 implies $z\in\rho(H_{N})$, hence $z$ is not an eigenvalue.
Assume that $K_{N}(z)$ is not invertible and take $0\neq g\in\text{\rm Ker}K_{N}(z)$. Set
$u:=-\Gamma(z)\Theta g.$
As in the proof of Theorem 5, we have $u\in H^{1}(\mathbb{R}^{3})$. Moreover, by Lemma 4(b),
$\tau u=-m(z)\Theta g,$
and $K_{N}(z)g=0$ is equivalent to $g=-m(z)\Theta g$, hence $\tau u=g$. Repeating the form computation in the proof of Theorem 5 with $f=0$ gives
$h_{N}[u,\varphi]-z(u,\varphi)_{L^{2}(\mathbb{R}^{3})}=0\qquad(\forall\,\varphi\in H^{1}(\mathbb{R}^{3})).$
Thus $u\in D(H_{N})$ and $(H_{N}-z)u=0$, so $z$ is an eigenvalue.
Conversely, let $0\neq u\in\text{\rm Ker}(H_{N}-z)$ and set $g:=\tau u$. Arguing as in the proof of Theorem 5 with $f=0$ (i.e., using the identity (3.7) for the resolvent difference) shows that $u=-\Gamma(z)\Theta g$ and $K_{N}(z)g=0$, hence $K_{N}(z)$ is not invertible; cf. Lemma 4 and the form characterization (3.6).
Finally, the maps $g\mapsto u=-\Gamma(z)\Theta g$ and $u\mapsto\tau u$ restrict to inverse bijections between $\text{\rm Ker}K_{N}(z)$ and $\text{\rm Ker}(H_{N}-z)$, and therefore $\dim\text{\rm Ker}(H_{N}-z)=\dim\text{\rm Ker}K_{N}(z)$. ∎
Before turning to rotational symmetry, we note that trace class perturbations of the resolvent are closely related to scattering theory. In the present setting Theorem 5 shows that $(H_{N}-z)^{-1}-R_{0}(z)$ is of trace class for all $z\in\rho(H_{N})$ with $z\in\mathbb{C}\setminus[0,\infty)$. By the Birman–Kuroda theorem [23], this implies existence and completeness of the wave operators and unitary equivalence of the absolutely continuous parts. As a consequence, the absolutely continuous and essential spectra of $H_{N}$ can be identified explicitly, as stated in the following theorem.
Theorem 8.
The absolutely continuous and essential spectra of $H_{N}$ coincide with those of the free Hamiltonian:
$\sigma_{\mathrm{ac}}(H_{N})=[0,\infty),\qquad\sigma_{\mathrm{ess}}(H_{N})=[0,\infty).$
Moreover, the negative spectrum of $H_{N}$ is purely discrete and consists of eigenvalues of finite multiplicity; the only possible accumulation point of the negative eigenvalues is 0.
Remark 2.
The operator $K_{N}(z)$ acts on an infinite–dimensional Hilbert space, so in general a global determinant is not defined. The correct spectral condition is the non–invertibility of $K_{N}(z)$. If we assume rotational symmetry, namely that each $\alpha_{j}$ is a constant, then $\Theta$ acts only on the shell index and the kernels of $m_{ij}(z)$ are rotation invariant. In that case one can reduce $K_{N}(z)$ by the spherical harmonic decomposition and obtain finite–dimensional secular equations in each angular momentum channel.
In the remainder of the paper, when we discuss explicit closed formulas and partial–wave reduction, we assume rotational symmetry, that is, $\alpha_{j}$ are constants.
Remark 3.
Under rotational symmetry the $\ell$–channel block $m_{\ell}(z)$ admits an explicit closed form in terms of spherical Bessel and Hankel functions. For the reader’s convenience we record the formula in Appendix B, together with the examples $N=2$ and $\ell=0,1$.
Lemma 9.
Let $\{Y_{\ell m}:-\ell\leq m\leq\ell\}$ be a complete orthonormal basis of spherical harmonics on $S^{2}$, and set
$\mathcal{K}_{N}=\bigoplus_{j=1}^{N}L^{2}(S^{2})=\bigoplus_{\ell=0}^{\infty}\mathcal{H}_{\ell}^{\oplus N},\qquad\mathcal{H}_{\ell}=\operatorname{span}\{Y_{\ell m}:-\ell\leq m\leq\ell\}.$
Suppose that the entries of $m(z)$ have rotation–invariant kernels $k_{ij}(\omega\cdot\omega^{\prime})$, $1\leq i,j\leq N$, and that $\Theta$ acts only on the shell index. Then for each $\ell\geq 0$ there is an $N\times N$ matrix $m_{\ell}(z)$ such that
$m(z)\big|_{\mathcal{H}_{\ell}^{\oplus N}}=m_{\ell}(z)\otimes I_{\mathcal{H}_{\ell}}.$
In particular, $K_{N}(z)$ fails to be invertible if and only if
$\det\bigl(I_{N}+m_{\ell}(z)\Theta\bigr)=0$
for some $\ell\geq 0$, and the corresponding eigenspace has dimension $(2\ell+1)\,\dim\text{\rm Ker}\bigl(I_{N}+m_{\ell}(z)\Theta\bigr)$.
Proof.
The spherical harmonic addition theorem (see, e.g., [1]) states that
$P_{\ell}(\omega\cdot\omega^{\prime})=\frac{4\pi}{2\ell+1}\sum_{m=-\ell}^{\ell}Y_{\ell m}(\omega)\,\overline{Y_{\ell m}(\omega^{\prime})},\qquad\omega,\omega^{\prime}\in S^{2},$
where $P_{\ell}$ is the Legendre polynomial of degree $\ell$. If we expand each rotation–invariant kernel $k_{ij}$ in Legendre polynomials and use the addition theorem, then $m(z)$ is diagonal in the spherical harmonic basis and does not mix different values of $\ell$ or $m$. More precisely, $m(z)$ leaves $\mathcal{H}_{\ell}^{\oplus N}$ invariant and acts there as
$m(z)\big|_{\mathcal{H}_{\ell}^{\oplus N}}=m_{\ell}(z)\otimes I_{\mathcal{H}_{\ell}},$
where $m_{\ell}(z)$ acts on the shell index.
Since $\Theta$ acts only on the shell index, it has the form $\Theta\otimes I_{\mathcal{H}_{\ell}}$ on $\mathcal{H}_{\ell}^{\oplus N}$, and hence
$K_{N}(z)\big|_{\mathcal{H}_{\ell}^{\oplus N}}=\bigl(I_{N}+m_{\ell}(z)\Theta\bigr)\otimes I_{\mathcal{H}_{\ell}}.$
Thus $K_{N}(z)$ is not invertible if and only if $I_{N}+m_{\ell}(z)\Theta$ is not invertible for some $\ell\geq 0$, that is, if and only if
$\det\bigl(I_{N}+m_{\ell}(z)\Theta\bigr)=0$
for some $\ell$. Moreover,
$\text{\rm Ker}\Bigl(\bigl(I_{N}+m_{\ell}(z)\Theta\bigr)\otimes I_{\mathcal{H}_{\ell}}\Bigr)=\text{\rm Ker}\bigl(I_{N}+m_{\ell}(z)\Theta\bigr)\otimes\mathcal{H}_{\ell},$
and therefore the stated dimension formula follows. ∎
Remark 4.
Equivalently, each $\mathcal{H}_{\ell}$ is an irreducible representation of ${\rm SO}(3)$, and every rotation–invariant operator acts as a scalar multiple of the identity on $\mathcal{H}_{\ell}$. In this language the NN–shell boundary space $\mathcal{H}_{\ell}^{\oplus N}$ is a direct sum of $N$ equivalent irreducible representations. Here $\mathrm{Comm}(\cdot)$ denotes the commutant, that is, the algebra of all bounded operators commuting with the ${\rm SO}(3)$–action. By Schur’s lemma, the commutant becomes isomorphic to $M_{N}(\mathbb{C})$ acting on the shell index and tensored with the identity on the representation space. That is,
$\mathrm{Comm}\bigl(\mathcal{H}_{\ell}^{\oplus N}\bigr)=M_{N}(\mathbb{C})\otimes I_{\mathcal{H}_{\ell}}.$
4 Variational preliminaries and higher partial waves
From this section on we specialize to the rotationally symmetric two–shell setting. That is, we fix radii $0<R_{1}<R_{2}$ and take constant couplings $\alpha_{1},\alpha_{2}\in\mathbb{R}$. We write $H:=H_{2}$ for the corresponding double $\delta$–shell Hamiltonian constructed by the quadratic form method in Section 2. The purpose of this section is to show that higher angular momentum channels cannot produce more negative eigenvalues than the ss–wave channel, and that the ground state, if it exists, lies in the sector $\ell=0$.
Since the interaction is radial, $H$ commutes with rotations and admits a partial–wave decomposition. We first recall a standard variational comparison lemma and then apply it to the radial quadratic forms.
Lemma 10.
Let $\mathcal{X}$ be a Hilbert space and let $T_{0}$ and $T$ be self–adjoint operators in $\mathcal{X}$ which are bounded from below and have the same closed form domain $Q$. Assume that the spectrum of each operator below its essential spectrum is discrete so that $\{\lambda_{n}(\cdot)\}_{n\geq 1}$ is well-defined. Denote their quadratic forms by $q_{0}$ and $q$. Assume that
$q[f]\geq q_{0}[f],\qquad f\in Q.$
Then the eigenvalues of $T$ and $T_{0}$, counted with multiplicity and listed in nondecreasing order, satisfy
$\lambda_{n}(T)\geq\lambda_{n}(T_{0})\qquad\text{for all }n\geq 1,$
and in particular
$N_{-}(T)\leq N_{-}(T_{0}),$
where $N_{-}(S)$ denotes the number of negative eigenvalues of a self–adjoint operator $S$.
Proof.
By the min–max principle (see e.g., [17]) one has
$\lambda_{n}(T)=\sup_{\begin{subarray}{c}L\subset Q\\ \dim L=n-1\end{subarray}}\ \inf_{\begin{subarray}{c}f\in Q\\ f\perp L,\,\|f\|=1\end{subarray}}q[f],\qquad\lambda_{n}(T_{0})=\sup_{\begin{subarray}{c}L\subset Q\\ \dim L=n-1\end{subarray}}\ \inf_{\begin{subarray}{c}f\in Q\\ f\perp L,\,\|f\|=1\end{subarray}}q_{0}[f].$
Since $q\geq q_{0}$ on $Q$, the inner infimum in the min–max formula satisfies
$\inf_{\begin{subarray}{c}f\in Q\\ f\perp L,\,\|f\|=1\end{subarray}}q[f]\geq\inf_{\begin{subarray}{c}f\in Q\\ f\perp L,\,\|f\|=1\end{subarray}}q_{0}[f]$
for every subspace $L\subset Q$ with $\dim L=n-1$. Taking the supremum over all such $L$ gives $\lambda_{n}(T)\geq\lambda_{n}(T_{0})$.
To compare the numbers of negative eigenvalues, let $N:=N_{-}(T_{0})$. If $N=\infty$, then the claim is trivial. Otherwise $\lambda_{N+1}(T_{0})\geq 0$ by definition of $N_{-}(T_{0})$, and hence
$\lambda_{N+1}(T)\geq\lambda_{N+1}(T_{0})\geq 0.$
Therefore $T$ has at most $N$ negative eigenvalues, that is, $N_{-}(T)\leq N_{-}(T_{0})$. ∎
We recall the standard partial–wave decomposition of $L^{2}(\mathbb{R}^{3})$ with respect to the action of the rotation group, under which the Hamiltonian $H$ reduces to a direct sum of one–dimensional radial operators labeled by the angular momentum $\ell$.
Theorem 11.
Fix radii $0<R_{1}<R_{2}$ and couplings $\alpha_{1},\alpha_{2}\in\mathbb{R}$, and let $H$ be the double $\delta$–shell Hamiltonian constructed in Section 2. Let $h_{\ell}$ denote the radial operator in the $\ell$–th partial wave. Then for every $\ell\geq 1$,
$\lambda_{k}(h_{\ell})\geq\lambda_{k}(h_{0})\qquad\text{for all }k\geq 1,$
and consequently
$N_{-}(h_{\ell})\leq N_{-}(h_{0}).$
Proof.
We recall the standard partial–wave decomposition (see, for example, [2, Sec. II.2]). Let $\{Y_{\ell m}\}_{\ell\geq 0,\,-\ell\leq m\leq\ell}$ be an orthonormal basis of spherical harmonics in $L^{2}(S^{2})$. Then $L^{2}(\mathbb{R}^{3})$ decomposes as the orthogonal sum
$L^{2}(\mathbb{R}^{3})=\bigoplus_{\ell=0}^{\infty}\bigoplus_{m=-\ell}^{\ell}\mathcal{H}_{\ell m},\qquad\mathcal{H}_{\ell m}=\Bigl\{u(x)=\frac{f(r)}{r}Y_{\ell m}(\omega):f\in L^{2}(0,\infty)\Bigr\},$
with $x=r\omega$, $r=|x|$, $\omega\in S^{2}$. Since the interaction is radial, $H$ commutes with rotations and leaves each $\mathcal{H}_{\ell m}$ invariant. On $\mathcal{H}_{\ell m}$ the operator $H$ is unitarily equivalent to a one–dimensional radial operator $h_{\ell}$ acting in $L^{2}(0,\infty)$, whose quadratic form is
$q_{\ell}[f]$ $=$ $\int_{0}^{\infty}\Bigl(|f^{\prime}(r)|^{2}+\frac{\ell(\ell+1)}{r^{2}}|f(r)|^{2}\Bigr)\,dr+\alpha_{1}|f(R_{1})|^{2}+\alpha_{2}|f(R_{2})|^{2},$
$\qquad f\in H^{1}(0,\infty)\ \text{with }f(0)=0.$
Here the point values $f(R_{j})$ are well defined for $f\in H^{1}(0,\infty)$. In particular, for each $\ell$ the negative eigenvalues of $H$ in the $\ell$–th partial wave coincide, with multiplicity, with those of $h_{\ell}$. Moreover, $h_{\ell}$ does not depend on $m$, so each eigenvalue of $h_{\ell}$ appears in $H$ with multiplicity $2\ell+1$.
For $\ell\geq 1$ the centrifugal term is nonnegative, and therefore
$q_{\ell}[f]=q_{0}[f]+\int_{0}^{\infty}\frac{\ell(\ell+1)}{r^{2}}|f(r)|^{2}\,dr\geq q_{0}[f],\qquad f\in H^{1}(0,\infty).$
Since $q_{\ell}\geq q_{0}$ on the common form domain $H^{1}(0,\infty)$ with $f(0)=0$, Lemma 10 implies
$\lambda_{k}(h_{\ell})\geq\lambda_{k}(h_{0})\qquad\text{for all }k\geq 1,$
and hence $N_{-}(h_{\ell})\leq N_{-}(h_{0})$. ∎
Remark 5.
Theorem 11 reduces the control of negative eigenvalues in all partial waves to the ss–wave channel $\ell=0$. Once the ss–wave problem is analyzed and $N_{-}(h_{0})$ is bounded, the same bound automatically holds for all $\ell\geq 1$.
As an immediate consequence, any negative eigenvalue of lowest energy must occur in the $\ell=0$ channel.
Corollary 12.
Under the assumptions of Theorem 11, if $H$ has at least one negative eigenvalue (cf. Theorem 5 and the trace class property of $(H-z)^{-1}-R_{0}(z)$), then the lowest eigenvalue belongs to the ss–wave sector $\ell=0$.
Proof.
Let $h_{\ell}$ be the radial operator in the $\ell$–th partial wave. For each $\ell\geq 1$ we have $q_{\ell}\geq q_{0}$ on $H^{1}(0,\infty)$ with $f(0)=0$, hence the min–max principle gives
$\lambda_{1}(h_{\ell})\geq\lambda_{1}(h_{0}).$
Therefore the smallest negative eigenvalue among all partial waves is attained in the ss–wave sector $\ell=0$, and the ground state of $H$ belongs to $\ell=0$. ∎
5 The two–shell case: detailed analysis of the ss–wave ($\ell=0$)
In the two–shell case $N=2$, the resolvent framework of Section 3 yields the spectral condition that the boundary operator $K(z)=I+m(z)\Theta$ fails to be invertible. On the other hand, the same eigenvalue problem admits a direct reduction to a one–dimensional radial ODE, leading to the usual coefficient matching conditions in each partial wave. The next lemma connects these two viewpoints: it reduces the operator condition to a finite–dimensional secular equation in every angular momentum channel and shows that this secular equation is equivalent to the radial matching condition.
Lemma 13.
Let $m(z)=(m_{ij}(z))_{i,j=1,2}$ be the boundary integral operator on $L^{2}(S^{2})\oplus L^{2}(S^{2})$ with kernel $G_{z}(R_{i}\omega,R_{j}\omega^{\prime})$, and let $\Theta=\mathrm{diag}(\alpha_{1}R_{1}^{2},\alpha_{2}R_{2}^{2})$. Assume $\text{{\rm Im}}\sqrt{z}>0$, that is, $z\in\mathbb{C}\setminus[0,\infty)$ on the principal branch. Then the following statements are equivalent.
(i) $z$ is an eigenvalue of $H$. Equivalently, $K(z)=I+m(z)\Theta$ is not invertible; cf. Proposition 7.
(ii) There exists $\ell\in\mathbb{N}\cup\{0\}$ such that
$\det\bigl(I+m_{\ell}(z)\Theta\bigr)=0,$
where $m_{\ell}(z)$ is the $2\times 2$ matrix describing the action of $m(z)$ on $\mathcal{H}_{\ell}\oplus\mathcal{H}_{\ell}$, that is,
$I+m_{\ell}(z)\Theta=\left(\begin{array}[]{cc}1+\alpha_{1}R_{1}^{2}m^{(11)}_{\ell}(z)&\alpha_{2}R_{2}^{2}m^{(12)}_{\ell}(z)\\[2.0pt] \alpha_{1}R_{1}^{2}m^{(21)}_{\ell}(z)&1+\alpha_{2}R_{2}^{2}m^{(22)}_{\ell}(z)\end{array}\right).$
Moreover, for each fixed $\ell$, the condition $\det\bigl(I+m_{\ell}(z)\Theta\bigr)=0$ is equivalent to the eigenvalue condition obtained from the corresponding radial problem in the $\ell$–channel.
Proof.
The kernel $G_{z}(R_{i}\omega,R_{j}\omega^{\prime})$ depends only on $\omega\cdot\omega^{\prime}$ and is therefore rotation invariant. Hence $m(z)$ commutes with the natural action of ${\rm SO}(3)$ on $L^{2}(S^{2})\oplus L^{2}(S^{2})$. Consequently, each spherical harmonic subspace
$\mathcal{H}_{\ell}:=\mathrm{span}\{Y_{\ell m}:\ -\ell\leq m\leq\ell\}$
is invariant, and $m(z)$ acts on $\mathcal{H}_{\ell}\oplus\mathcal{H}_{\ell}$ as a $2\times 2$ matrix $m_{\ell}(z)$ on the shell index. Equivalently, $m(z)$ is block diagonal with respect to
$L^{2}(S^{2})\oplus L^{2}(S^{2})=\bigoplus_{\ell=0}^{\infty}\bigoplus_{m=-\ell}^{\ell}\bigl(\mathbb{C}\,Y_{\ell m}\oplus\mathbb{C}\,Y_{\ell m}\bigr),$
and on each block it acts as
$m_{\ell}(z)=\left(\begin{array}[]{cc}m^{(11)}_{\ell}(z)&m^{(12)}_{\ell}(z)\\[2.0pt] m^{(21)}_{\ell}(z)&m^{(22)}_{\ell}(z)\end{array}\right).$
Therefore $K(z)=I+m(z)\Theta$ is unitarily equivalent to the orthogonal direct sum of the matrices $I+m_{\ell}(z)\Theta$, where each block appears with multiplicity $2\ell+1$. Since $m(z)$ is compact by Lemma 6, the operator $K(z)$ is Fredholm of index zero.
If $\det(I+m_{\ell}(z)\Theta)\neq 0$ for every $\ell$, then each block $I+m_{\ell}(z)\Theta$ is invertible and hence $\text{\rm Ker}K(z)=\{0\}$. For a Fredholm operator of index zero, injectivity implies surjectivity, so $K(z)$ is invertible. Conversely, if $\det(I+m_{\ell}(z)\Theta)=0$ for some $\ell$, then the corresponding block has a nontrivial kernel, and so does $K(z)$. Thus $K(z)$ is not invertible if and only if there exists $\ell$ such that $\det(I+m_{\ell}(z)\Theta)=0$, which proves the equivalence of (i) and (ii).
We now relate the condition $\det(I+m_{\ell}(z)\Theta)=0$ to the eigenvalue condition arising from the coefficient matching in the radial problem. Let $g={}^{t}(g_{1},g_{2})\neq 0$ satisfy $(I+m_{\ell}(z)\Theta)g=0$ and define
$u=-\sum_{j=1}^{2}\Gamma_{j}(z)\,(\Theta g)_{j},\qquad(\Theta g)_{j}=\alpha_{j}R_{j}^{2}g_{j}.$
Then $u$ solves $(-\Delta-z)u=0$ away from $S_{1}\cup S_{2}$. Since $g_{j}\in\mathcal{H}_{\ell}$, the angular dependence of $u$ lies in $\mathcal{H}_{\ell}$ and is proportional to $Y_{\ell m}$ for some $m$. We represent $u$ in the form
$u(x)=\frac{u_{\ell}(r)}{r}\,Y_{\ell m}(\omega),\qquad x=r\omega.$
The mapping properties of $\Gamma_{j}(z)$ then imply that on each interval $(0,R_{1})$, $(R_{1},R_{2})$, and $(R_{2},\infty)$ the function $u_{\ell}$ is a linear combination of the standard radial solutions of $(-\Delta-z)u=0$, expressed in terms of spherical Bessel and Hankel functions.
Taking traces on $S_{1}$ and $S_{2}$ and using the $\delta$–shell jump laws, we recall that $u$ is continuous at $r=R_{j}$ and satisfies
$\partial_{r}u(R_{j}+0,\omega)-\partial_{r}u(R_{j}-0,\omega)=\alpha_{j}u(R_{j},\omega),\qquad j=1,2.$
Moreover, by the trace identity $\tau_{i}\Gamma_{j}(z)=m_{ij}(z)$ and the standard jump relations for single–layer potentials (cf. Lemma 6 and the proof of Proposition 7), the resulting matching conditions yield a homogeneous linear system for the $\ell$–channel boundary data. By the definition of $m_{\ell}(z)$ (see Appendix B, in particular Lemma 18), this boundary system is exactly $(I+m_{\ell}(z)\Theta)g=0$. Thus a nontrivial vector $g$ in $\text{\rm Ker}(I+m_{\ell}(z)\Theta)$ produces a nontrivial radial solution satisfying the matching conditions, and conversely every such radial solution produces a nontrivial $g$. Hence $\det(I+m_{\ell}(z)\Theta)=0$ is equivalent to the radial matching secular equation in the $\ell$–channel. ∎
We now specialize to the ss–wave sector $\ell=0$ and derive an explicit eigenvalue condition by solving the corresponding radial problem.
Proposition 14.
Let $0<R_{1}<R_{2}$, $\alpha_{1},\alpha_{2}\in\mathbb{R}$, and set $d=R_{2}-R_{1}$. In the ss–wave sector ($\ell=0$) we write $u(r)=rf(r)$. Then $u$ satisfies
$u^{\prime\prime}(r)=E\,u(r)\quad\text{for }r\neq R_{1},R_{2},\qquad u(0)=0,\qquad u\in L^{2}(0,\infty),$ (5.1)
together with the interface conditions
[…]
For $E=-\kappa^{2}<0$ the ss–wave eigenvalues are precisely those numbers $E=-\kappa^{2}$ with $\kappa>0$ which satisfy
$F_{d}(\kappa):=\Bigl[\gamma_{1}(\kappa)+(\alpha_{2}+\kappa)\Bigr]\cosh(\kappa d)+\Bigl[\kappa+\gamma_{1}(\kappa)+\frac{\alpha_{2}\,\gamma_{1}(\kappa)}{\kappa}\Bigr]\sinh(\kappa d)=0,$ (5.3)
where
$\gamma_{1}(\kappa)=\alpha_{1}+\kappa\,\coth(\kappa R_{1}).$ (5.4)
Equivalently, nontrivial solutions exist if and only if $\det M(\kappa;d)=0$, where
$M(\kappa;d)=\left(\begin{array}[]{cc}-\gamma_{1}(\kappa)&\kappa\\[4.0pt] (\alpha_{2}+\kappa)\cosh(\kappa d)+\kappa\sinh(\kappa d)&(\alpha_{2}+\kappa)\sinh(\kappa d)+\kappa\cosh(\kappa d)\end{array}\right).$ (5.5)
Proof.
We take $E=-\kappa^{2}<0$ with $\kappa>0$. The general solution of (5.1) in the three regions $0<r<R_{1}$, $R_{1}<r<R_{2}$, and $r>R_{2}$ can be written as
$u(r)=\begin{cases}A\,\sinh(\kappa r),&0<r<R_{1},\\[4.30554pt] B\,e^{\kappa r}+C\,e^{-\kappa r},&R_{1}<r<R_{2},\\[4.30554pt] D\,e^{-\kappa r},&r>R_{2},\end{cases}$ (5.6)
with constants $A,B,C,D\in\mathbb{C}$. Continuity at $R_{1}$ and $R_{2}$ gives
$A\sinh(\kappa R_{1})$ $=$ $Be^{\kappa R_{1}}+Ce^{-\kappa R_{1}},$ (5.7)
$De^{-\kappa R_{2}}$ $=$ $Be^{\kappa R_{2}}+Ce^{-\kappa R_{2}}.$ (5.8)
Differentiating (5.6) yields
$u_{1}^{\prime}(R_{1})$ $=$ $A\kappa\cosh(\kappa R_{1}),$
$u_{2}^{\prime}(R_{1})$ $=$ $\kappa\bigl(Be^{\kappa R_{1}}-Ce^{-\kappa R_{1}}\bigr),$
$u_{2}^{\prime}(R_{2})$ $=$ $\kappa\bigl(Be^{\kappa R_{2}}-Ce^{-\kappa R_{2}}\bigr),$
$u_{3}^{\prime}(R_{2})$ $=$ $-\,\kappa De^{-\kappa R_{2}}.$ (5.9)
Using the jump conditions (5.2) at $R_{1}$ and $R_{2}$ we obtain
$\kappa\bigl(Be^{\kappa R_{1}}-Ce^{-\kappa R_{1}}\bigr)-A\kappa\cosh(\kappa R_{1})$ $=$ $\alpha_{1}\bigl(Be^{\kappa R_{1}}+Ce^{-\kappa R_{1}}\bigr),$ (5.10)
$-\kappa De^{-\kappa R_{2}}-\kappa\bigl(Be^{\kappa R_{2}}-Ce^{-\kappa R_{2}}\bigr)$ $=$ $\alpha_{2}\bigl(Be^{\kappa R_{2}}+Ce^{-\kappa R_{2}}\bigr).$ (5.11)
From (5.7) we eliminate $A$ and $D$ and obtain
$A=\frac{Be^{\kappa R_{1}}+Ce^{-\kappa R_{1}}}{\sinh(\kappa R_{1})},\qquad De^{-\kappa R_{2}}=Be^{\kappa R_{2}}+Ce^{-\kappa R_{2}}.$ (5.12)
Substituting (5.12) into (5.10) yields
$\kappa\bigl(Be^{\kappa R_{1}}-Ce^{-\kappa R_{1}}\bigr)$ $=$ $\bigl(\alpha_{1}+\kappa\coth(\kappa R_{1})\bigr)\bigl(Be^{\kappa R_{1}}+Ce^{-\kappa R_{1}}\bigr),$ (5.13)
$\kappa\bigl(Be^{\kappa R_{2}}-Ce^{-\kappa R_{2}}\bigr)$ $=$ $-\bigl(\alpha_{2}+\kappa\bigr)\bigl(Be^{\kappa R_{2}}+Ce^{-\kappa R_{2}}\bigr).$ (5.14)
We now introduce the combinations
$X(r)=Be^{\kappa r}+Ce^{-\kappa r},\qquad Y(r)=Be^{\kappa r}-Ce^{-\kappa r}.$ (5.15)
Then (5.13) can be rewritten as
$\kappa Y(R_{1})=\gamma_{1}(\kappa)X(R_{1}),\qquad\kappa Y(R_{2})=-(\alpha_{2}+\kappa)X(R_{2}),$ (5.16)
where $\gamma_{1}(\kappa)$ is given by (5.4). Between $R_{1}$ and $R_{2}$ the pair $(X,Y)$ satisfies
$X(R_{2})$ $=$ $\cosh(\kappa d)\,X(R_{1})+\sinh(\kappa d)\,Y(R_{1}),$ (5.17)
$Y(R_{2})$ $=$ $\sinh(\kappa d)\,X(R_{1})+\cosh(\kappa d)\,Y(R_{1}).$ (5.18)
From (5.16) at $R_{1}$ we have
$Y(R_{1})=\frac{\gamma_{1}(\kappa)}{\kappa}\,X(R_{1}).$ (5.19)
Substituting (5.19) into (5.17) and (5.18) yields
$X(R_{2})$ $=$ $\Bigl(\cosh(\kappa d)+\frac{\gamma_{1}(\kappa)}{\kappa}\sinh(\kappa d)\Bigr)X(R_{1}),$ (5.20)
$Y(R_{2})$ $=$ $\Bigl(\sinh(\kappa d)+\frac{\gamma_{1}(\kappa)}{\kappa}\cosh(\kappa d)\Bigr)X(R_{1}).$ (5.21)
Applying the boundary condition at $R_{2}$ in (5.16), we obtain a homogeneous linear system for $X(R_{1})$ and $Y(R_{1})$,
$M(\kappa;d)\left(\begin{array}[]{c}X(R_{1})\\[2.0pt] Y(R_{1})\end{array}\right)=0,$
with $M(\kappa;d)$ as in (5.5). There exists a nontrivial solution if and only if $\det M(\kappa;d)=0$. A direct computation of the determinant gives (5.3), and the statement follows. ∎
6 Counting of ss–wave bound states for large shell separation
In this section we restrict to the ss–wave sector $\ell=0$ and study how the number of negative eigenvalues behaves when the shell separation $d=R_{2}-R_{1}$ becomes large. When the two $\delta$–shells are far apart, bound states originating from attractive single shells interact only weakly and the coupling is exponentially small in $d$. As a result, the number of negative ss–wave eigenvalues stabilizes and can be read off from the corresponding one–shell problems.
Proposition 15.
Let $H$ be the double $\delta$–shell Hamiltonian with radii $0<R_{1}<R_{2}$ and couplings $\alpha_{1},\alpha_{2}\in\mathbb{R}$. In the ss–wave subspace, the negative eigenvalues $E=-\kappa^{2}$ are in one–to–one correspondence with the positive roots $\kappa>0$ of the secular equation (5.3).
The associated ss–wave quadratic form is given by
$q_{0}[f]=\int_{0}^{\infty}|f^{\prime}(r)|^{2}\,dr+\alpha_{1}|f(R_{1})|^{2}+\alpha_{2}|f(R_{2})|^{2},\qquad f\in H^{1}(0,\infty).$
Hence the negative part of $q_{0}$ depends only on the two point evaluations $f\mapsto f(R_{1})$ and $f\mapsto f(R_{2})$. As a consequence, the ss–wave radial operator can have at most two negative eigenvalues (counted with multiplicity), and therefore the secular equation (5.3) admits at most two positive roots.
Proof.
By Proposition 14 the negative ss–wave eigenvalues $E=-\kappa^{2}$ are characterized by the positive roots $\kappa>0$ of $F_{d}(\kappa)=0$, where $F_{d}$ is given in (5.3). We therefore study the function $F_{d}(\kappa)$.
The ss–wave secular equation has the form
$F_{d}(\kappa)=A(\kappa)\cosh(\kappa d)+B(\kappa)\sinh(\kappa d),\qquad\kappa>0,$
where
$A(\kappa)$ $=$ $\gamma_{1}(\kappa)+\alpha_{2}+\kappa,$
$B(\kappa)$ $=$ $\kappa+\gamma_{1}(\kappa)+\frac{\alpha_{2}\gamma_{1}(\kappa)}{\kappa},$
$\gamma_{1}(\kappa)$ $=$ $\alpha_{1}+\kappa\coth(\kappa R_{1}).$
Using $\cosh t=\tfrac{1}{2}(e^{t}+e^{-t})$ and $\sinh t=\tfrac{1}{2}(e^{t}-e^{-t})$, we obtain
$F_{d}(\kappa)=\tfrac{1}{2}\bigl[F_{\infty}(\kappa)e^{\kappa d}+G(\kappa)e^{-\kappa d}\bigr],$ (6.1)
where
$F_{\infty}(\kappa)=\frac{(\kappa+\gamma_{1}(\kappa))(2\kappa+\alpha_{2})}{\kappa},\qquad G(\kappa)=\alpha_{2}\Bigl(1-\frac{\gamma_{1}(\kappa)}{\kappa}\Bigr).$ (6.2)
We first show that the ss–wave radial operator has at most two negative eigenvalues. Assume for contradiction that it has at least three negative eigenvalues and let $L$ be the span of three corresponding eigenfunctions. Then $\dim L=3$ and
$q_{0}[f]=(h_{0}f,f)_{L^{2}(0,\infty)}<0\qquad\text{for all }0\neq f\in L.$
The linear map $f\mapsto(f(R_{1}),f(R_{2}))\in\mathbb{C}^{2}$ cannot be injective on $L$, hence there exists $0\neq f\in L$ such that $f(R_{1})=f(R_{2})=0$. For this $f$ we have
$q_{0}[f]=\int_{0}^{\infty}|f^{\prime}(r)|^{2}\,dr>0,$
since $\int_{0}^{\infty}|f^{\prime}(r)|^{2}\,dr=0$ would imply that $f$ is constant, and the conditions $f(R_{1})=f(R_{2})=0$ would force $f\equiv 0$. This contradiction shows $N_{-}(h_{0})\leq 2$, and therefore $F_{d}(\kappa)=0$ has at most two positive roots, counted with multiplicity.
(i) We first consider the case $\alpha_{1},\alpha_{2}\geq 0$. Then $\gamma_{1}(\kappa)\geq 0$, $A(\kappa)>0$, and $B(\kappa)\geq 0$ for all $\kappa>0$. Since $\cosh(\kappa d)>0$ and $\sinh(\kappa d)>0$ for all $\kappa>0$ and $d>0$, we have $F_{d}(\kappa)>0$ for all $\kappa>0$. Hence, there is no positive root of $F_{d}$, and therefore no negative ss–wave eigenvalue.
(ii) Next we assume that exactly one shell is attractive and that the corresponding one–shell problem supports an ss–wave bound state.
Case 1: $\alpha_{1}<0\leq\alpha_{2}$ and $\alpha_{1}<-1/R_{1}$. Then, there exists a unique $\kappa_{\mathrm{in}}>0$ satisfying
$\kappa+\gamma_{1}(\kappa)=0,\qquad\text{that is,}\qquad\alpha_{1}+\kappa\coth(\kappa R_{1})+\kappa=0.$
At $\kappa=\kappa_{\mathrm{in}}$ one has $F_{\infty}(\kappa_{\mathrm{in}})=0$. Moreover $F_{\infty}^{\prime}(\kappa_{\mathrm{in}})\neq 0$. Indeed, $2\kappa_{\mathrm{in}}+\alpha_{2}>0$ and $\kappa_{\mathrm{in}}+\gamma_{1}(\kappa_{\mathrm{in}})=0$ imply $F_{\infty}^{\prime}(\kappa_{\mathrm{in}})>0$. Expanding (6.1) near $\kappa_{\mathrm{in}}$ yields
$\kappa(d)=\kappa_{\mathrm{in}}-\frac{G(\kappa_{\mathrm{in}})}{F^{\prime}_{\infty}(\kappa_{\mathrm{in}})}e^{-2\kappa_{\mathrm{in}}d}+O(e^{-4\kappa_{\mathrm{in}}d}),$
so for all sufficiently large $d$ there exists exactly one positive root of $F_{d}$. The corresponding eigenvalue gives the unique negative ss–wave eigenvalue in this case.
Case 2: $\alpha_{2}<0\leq\alpha_{1}$ and $\alpha_{2}<-1/R_{2}$. For the single outer shell at $r=R_{2}$, the ss–wave bound state condition is $\alpha_{2}<-1/R_{2}$, and the corresponding root is given by the unique positive zero of $2\kappa+\alpha_{2}$, namely
$\kappa_{\mathrm{out}}=-\frac{\alpha_{2}}{2}>0.$
At $\kappa=\kappa_{\mathrm{out}}$ we have $F_{\infty}(\kappa_{\mathrm{out}})=0$. Moreover
$F_{\infty}^{\prime}(\kappa_{\mathrm{out}})=\frac{2\bigl(\kappa_{\mathrm{out}}+\gamma_{1}(\kappa_{\mathrm{out}})\bigr)}{\kappa_{\mathrm{out}}}\neq 0,$
and since $\alpha_{1}\geq 0$ implies $\gamma_{1}(\kappa_{\mathrm{out}})>0$, we have $F_{\infty}^{\prime}(\kappa_{\mathrm{out}})>0$. Expanding (6.1) near $\kappa_{\mathrm{out}}$ yields
$\kappa(d)=\kappa_{\mathrm{out}}-\frac{G(\kappa_{\mathrm{out}})}{F^{\prime}_{\infty}(\kappa_{\mathrm{out}})}e^{-2\kappa_{\mathrm{out}}d}+O(e^{-4\kappa_{\mathrm{out}}d}).$
Thus, we again obtain exactly one positive root for all sufficiently large $d$. This root gives the unique negative ss–wave eigenvalue in this case.
(iii) Finally we consider the case when both shells are attractive, $\alpha_{1}<0$ and $\alpha_{2}<0$. Assume that each one–shell subsystem supports an ss–wave bound state, that is, $\alpha_{j}<-1/R_{j}$ for $j=1,2$. Let $\kappa_{\mathrm{in}},\kappa_{\mathrm{out}}>0$ be the corresponding isolated roots, that is, the zeros of $\kappa+\gamma_{1}(\kappa)$ and $2\kappa+\alpha_{2}$, respectively. Then
$F_{\infty}(\kappa_{\mathrm{in}})=0,\qquad F_{\infty}(\kappa_{\mathrm{out}})=0.$
Assume that these zeros are simple, that is, $F_{\infty}^{\prime}(\kappa_{\mathrm{in}})\neq 0$ and $F_{\infty}^{\prime}(\kappa_{\mathrm{out}})\neq 0$. (If one of them is not simple, then the corresponding critical tuning is treated separately in the tunneling regime.) For large $d$ the roots are only slightly perturbed, and we can write
$\delta_{\mathrm{in}}(d)$ $=$ $-\frac{G(\kappa_{\mathrm{in}})}{F^{\prime}_{\infty}(\kappa_{\mathrm{in}})}e^{-2\kappa_{\mathrm{in}}d}+O(e^{-4\kappa_{\mathrm{in}}d}),$
$\delta_{\mathrm{out}}(d)$ $=$ $-\frac{G(\kappa_{\mathrm{out}})}{F^{\prime}_{\infty}(\kappa_{\mathrm{out}})}e^{-2\kappa_{\mathrm{out}}d}+O(e^{-4\kappa_{\mathrm{out}}d}),$
so for all sufficiently large $d$ there are two distinct positive roots of $F_{d}$, namely $\kappa_{\mathrm{in}}+\delta_{\mathrm{in}}(d)$ and $\kappa_{\mathrm{out}}+\delta_{\mathrm{out}}(d)$.
For general values of $d$, the number of positive roots in $(0,\infty)$ can change only when $F_{d}$ has a multiple root $\kappa_{0}>0$, that is,
$\begin{cases}A(\kappa_{0})\cosh(\kappa_{0}d)+B(\kappa_{0})\sinh(\kappa_{0}d)=0,\\ (A^{\prime}(\kappa_{0})+dB(\kappa_{0}))\cosh(\kappa_{0}d)+(B^{\prime}(\kappa_{0})+dA(\kappa_{0}))\sinh(\kappa_{0}d)=0.\end{cases}$ (6.3)
We exclude here the threshold $\kappa=0$. This gives a codimension–one condition on the parameters $(\alpha_{1},\alpha_{2},d)$. In the generic situation the two roots persist and merge only at isolated critical values of $(\alpha_{1},\alpha_{2},d)$. ∎
Corollary 16.
Assume that exactly one of $\alpha_{1},\alpha_{2}$ is negative and that the corresponding single $\delta$–shell does not support an ss–wave bound state, that is,
$\alpha_{1}\geq-\frac{1}{R_{1}}\ \text{if }\alpha_{2}\geq 0,\qquad\text{or}\qquad\alpha_{2}\geq-\frac{1}{R_{2}}\ \text{if }\alpha_{1}\geq 0.$
Then the ss–wave secular equation (5.3) has no positive root. Consequently, $H$ has no negative eigenvalue.
Proof.
We first show that the ss–wave secular equation has no positive root.
Case 1: $\alpha_{1}<0\leq\alpha_{2}$ and $\alpha_{1}\geq-1/R_{1}$. For $\kappa>0$ one has $\kappa\coth(\kappa R_{1})\geq 1/R_{1}$, hence $\gamma_{1}(\kappa)=\alpha_{1}+\kappa\coth(\kappa R_{1})\geq 0$. Therefore
$A(\kappa)=\gamma_{1}(\kappa)+\alpha_{2}+\kappa>0,\qquad B(\kappa)=\kappa+\gamma_{1}(\kappa)+\frac{\alpha_{2}\gamma_{1}(\kappa)}{\kappa}\geq 0,$
and since $\cosh(\kappa d)>0$ and $\sinh(\kappa d)>0$ for $d>0$ we obtain $F_{d}(\kappa)>0$ for all $\kappa>0$. Thus (5.3) has no positive root.
Case 2: $\alpha_{2}<0\leq\alpha_{1}$ and $\alpha_{2}\geq-1/R_{2}$. Let $q_{0}$ be the ss–wave quadratic form,
$q_{0}[f]=\int_{0}^{\infty}|f^{\prime}(r)|^{2}\,dr+\alpha_{1}|f(R_{1})|^{2}+\alpha_{2}|f(R_{2})|^{2},\qquad f\in H^{1}(0,\infty).$
Since $\alpha_{1}\geq 0$, we have
$q_{0}[f]\geq\int_{0}^{\infty}|f^{\prime}(r)|^{2}\,dr+\alpha_{2}|f(R_{2})|^{2}=:q_{\mathrm{out}}[f].$
The form $q_{\mathrm{out}}$ is the ss–wave form for the single outer shell at $r=R_{2}$, and the hypothesis $\alpha_{2}\geq-1/R_{2}$ means that this one–shell problem has no negative ss–wave eigenvalue. By Lemma 10 it follows that the double–shell ss–wave operator $h_{0}$ also has no negative eigenvalue. Hence (5.3) has no positive root in this case.
Having excluded negative eigenvalues in the ss–wave channel, we treat higher partial waves. For $\ell\geq 1$ the radial quadratic form is
$q_{\ell}[f]=\int_{0}^{\infty}|f^{\prime}(r)|^{2}\,dr+\ell(\ell+1)\int_{0}^{\infty}\frac{|f(r)|^{2}}{r^{2}}\,dr+\alpha_{1}|f(R_{1})|^{2}+\alpha_{2}|f(R_{2})|^{2},$
so $q_{\ell}[f]\geq q_{0}[f]$ for all $f\in H^{1}(0,\infty)$. Another application of Lemma 10 gives $N_{-}(h_{\ell})\leq N_{-}(h_{0})=0$ for every $\ell\geq 1$. Therefore $H$ has no negative eigenvalue. ∎
Remark 6.
In the outer–attractive case $\alpha_{2}<0\leq\alpha_{1}$, the one–shell threshold $1/R_{2}$ tends to $0^{-}$ as $R_{2}=R_{1}+d\to\infty$. Hence any fixed $\alpha_{2}<0$ eventually satisfies $\alpha_{2}<-1/R_{2}$, and for all sufficiently large $d$ the ss–wave secular equation acquires a positive root and produces a negative ss–wave eigenvalue as in Proposition 15(ii).
7 Tunneling Splitting of Eigenvalues
In this section we analyze the tunneling splitting of the lowest ss–wave eigenvalues for the double $\delta$–shell operator in the regime of large shell separation. As shown in Section 6, when the two shells are well separated and the corresponding one–shell eigenvalues are distinct, each ss–wave eigenvalue converges to its one–shell counterpart with an exponentially small correction of order $e^{-2\kappa d}$.
To isolate a genuine tunneling effect, we tune the parameters so that the limiting one–shell eigenvalues coincide at a common energy $E_{0}=-\kappa_{0}^{2}<0$. In this critical situation, the standard perturbative picture breaks down and the interaction between the two shells produces a pair of eigenvalues whose splitting occurs on the larger scale $e^{-\kappa_{0}d}$. This regime is analogous to the symmetric double–well situation in one–dimensional quantum mechanics and provides a natural setting in which to compare the exact splitting with the prediction based on Agmon distances. We now establish this splitting explicitly within the present solvable model.
Theorem 17 (Tunneling splitting of ss–wave eigenvalues).
Let $R_{1}>0$ and $\alpha_{1},\alpha_{2}\in\mathbb{R}$. Assume that there exists $\kappa_{0}>0$ such that
$\gamma_{1}(\kappa_{0})+\kappa_{0}=0,\qquad\alpha_{2}+2\kappa_{0}=0,$ (7.1)
where $\gamma_{1}(\kappa)=\alpha_{1}+\kappa\coth(\kappa R_{1})$. Then, as $d=R_{2}-R_{1}\to\infty$, the secular equation admits two solutions $\kappa_{\pm}(d)$ near $\kappa_{0}$, and they satisfy
$\kappa_{\pm}(d)=\kappa_{0}\pm C\,e^{-\kappa_{0}d}+o(e^{-\kappa_{0}d}),$ (7.2)
with the constant
$C=\left(-\,\frac{2\,G(\kappa_{0})}{F_{\infty}^{\prime\prime}(\kappa_{0})}\right)^{1/2}>0.$ (7.3)
Consequently,
$E_{\pm}(d)=-\kappa_{\pm}(d)^{2},\qquad\bigl|E_{+}(d)-E_{-}(d)\bigr|=4\kappa_{0}C\,e^{-\kappa_{0}d}+o(e^{-\kappa_{0}d}).$ (7.4)
Proof.
We use the representation (6.1),
$F_{d}(\kappa)=\tfrac{1}{2}\bigl[F_{\infty}(\kappa)e^{\kappa d}+G(\kappa)e^{-\kappa d}\bigr],\qquad F_{\infty}(\kappa)=\frac{(\kappa+\gamma_{1}(\kappa))(2\kappa+\alpha_{2})}{\kappa},$
where $G$ is given in (6.2). The tuning conditions (7.1) imply
$\kappa_{0}+\gamma_{1}(\kappa_{0})=0,\qquad 2\kappa_{0}+\alpha_{2}=0,$
and therefore $F_{\infty}(\kappa_{0})=0$.
Differentiating $F_{\infty}$ and using the fact that both factors vanish at $\kappa_{0}$, we obtain $F_{\infty}^{\prime}(\kappa_{0})=0$, so $\kappa_{0}$ is at least a double zero of $F_{\infty}$. Moreover,
$\gamma_{1}^{\prime}(\kappa)=\coth(\kappa R_{1})-\kappa R_{1}\,\mathrm{csch}^{2}(\kappa R_{1})=\frac{\sinh(2t)-2t}{2\sinh^{2}t}>0,\qquad t=\kappa R_{1}>0,$
and hence
$F_{\infty}^{\prime\prime}(\kappa_{0})=\frac{4\,(1+\gamma_{1}^{\prime}(\kappa_{0}))}{\kappa_{0}}>0.$
Furthermore, by (6.2) and $2\kappa_{0}+\alpha_{2}=0$ we have
$G(\kappa_{0})=\alpha_{2}\Bigl(1-\frac{\gamma_{1}(\kappa_{0})}{\kappa_{0}}\Bigr)=\alpha_{2}\Bigl(1-(-1)\Bigr)=2\alpha_{2}=-4\kappa_{0}<0.$
Therefore
$\frac{2\,G(\kappa_{0})}{F_{\infty}^{\prime\prime}(\kappa_{0})}>0,$
and the constant $C$ in (7.3) is well defined and strictly positive. Since $F_{\infty}(\kappa_{0})=F_{\infty}^{\prime}(\kappa_{0})=0$ and $F_{\infty}^{\prime\prime}(\kappa_{0})>0$, the zero of $F_{\infty}$ at $\kappa_{0}$ has multiplicity exactly two.
Existence of two roots near $\kappa_{0}$ for large $d$. Fix $\delta_{0}\in(0,\kappa_{0}/2)$ so that $F_{\infty}$ and $G$ are $C^{3}$ on $\kappa_{0}-\delta_{0},\kappa_{0}+\delta_{0}]$. Then the Taylor remainder
$F_{\infty}(\kappa_{0}+\delta)=\tfrac{1}{2}\,F_{\infty}^{\prime\prime}(\kappa_{0})\,\delta^{2}+O(\delta^{3})$ (7.5)
holds uniformly for $|\delta|\leq\delta_{0}$, and $G(\kappa)=G(\kappa_{0})+O(\delta)$ holds uniformly on the same interval.
Write $\kappa=\kappa_{0}+\delta$ and introduce the scaling
$\delta=\eta e^{-\kappa_{0}d},\qquad|\eta|\leq M,$
where $M>0$ is fixed. For all sufficiently large $d$ we have $|\delta|\leq\delta_{0}$. From (6.1) we can rewrite $F_{d}(\kappa)=0$ as
$F_{\infty}(\kappa)+G(\kappa)e^{-2\kappa d}=0.$ (7.6)
Multiplying (7.6) by $e^{2\kappa_{0}d}$ and using $\delta d=\eta de^{-\kappa_{0}d}=o(1)$ uniformly for $|\eta|\leq M$, we obtain
$e^{2\kappa_{0}d}F_{\infty}(\kappa_{0}+\eta e^{-\kappa_{0}d})+G(\kappa_{0})e^{-2\delta d}+o(1)=0\qquad(d\to\infty),$
where the $o(1)$ term is uniform for $|\eta|\leq M$. Using (7.5) with $\delta=\eta e^{-\kappa_{0}d}$ and $e^{-2\delta d}=1+o(1)$ uniformly for $|\eta|\leq M$, we arrive at
$H_{d}(\eta):=e^{2\kappa_{0}d}\,F_{\infty}(\kappa_{0}+\eta e^{-\kappa_{0}d})+G(\kappa_{0})=Q(\eta)+r_{d}(\eta),$ (7.7)
where
$Q(\eta):=\tfrac{1}{2}F_{\infty}^{\prime\prime}(\kappa_{0})\eta^{2}+G(\kappa_{0}),\qquad\sup_{|\eta|\leq M}|r_{d}(\eta)|\to 0\quad(d\to\infty).$
The polynomial $Q$ has exactly two simple real zeros $\eta=\pm C$, where $C=\bigl(-2G(\kappa_{0})/F_{\infty}^{\prime\prime}(\kappa_{0})\bigr)^{1/2}$. Choose $\varepsilon\in(0,C)$ so that $Q(C-\varepsilon)$ and $Q(C+\varepsilon)$ have opposite signs, and likewise $Q(-C-\varepsilon)$ and $Q(-C+\varepsilon)$ have opposite signs. By (7.7) and the uniform convergence $H_{d}\to Q$ on the two compact intervals $C-\varepsilon,C+\varepsilon]$ and $C-\varepsilon,-C+\varepsilon]$, the same sign pattern holds for $H_{d}$ for all sufficiently large $d$. Hence, by the intermediate value theorem, there exist
$\eta_{+}(d)\in(C-\varepsilon,C+\varepsilon),\qquad\eta_{-}(d)\in(-C-\varepsilon,-C+\varepsilon)$
such that $H_{d}(\eta_{\pm}(d))=0$.
To show $\eta_{\pm}(d)\to\pm C$, take any sequence $d_{n}\to\infty$. Since $\eta_{+}(d_{n})\in[C-\varepsilon,C+\varepsilon]$, there exists a subsequence $\eta_{+}(d_{n_{k}})\to\eta_{*}$ with $\eta_{*}\in[C-\varepsilon,C+\varepsilon]$. Because $H_{d_{n_{k}}}(\eta_{+}(d_{n_{k}}))=0$ and $H_{d_{n_{k}}}\to Q$ uniformly on $C-\varepsilon,C+\varepsilon]$, we obtain $Q(\eta_{*})=0$, hence $\eta_{*}=C$. Thus $\eta_{+}(d)\to C$, and similarly $\eta_{-}(d)\to-C$.
Define
$\kappa_{\pm}(d):=\kappa_{0}+\eta_{\pm}(d)e^{-\kappa_{0}d},\qquad\delta_{\pm}(d):=\kappa_{\pm}(d)-\kappa_{0}.$
Then $F_{d}(\kappa_{\pm}(d))=0$ and $\eta_{\pm}(d)=\pm C+o(1)$. Therefore
$\delta_{\pm}(d)=\pm\,C\,e^{-\kappa_{0}d}+o(e^{-\kappa_{0}d}),$
which proves (7.2).
Finally,
$E_{\pm}(d)=-\kappa_{\pm}(d)^{2}=-\bigl(\kappa_{0}+\delta_{\pm}(d)\bigr)^{2}=-\kappa_{0}^{2}-2\kappa_{0}\,\delta_{\pm}(d)+O(\delta_{\pm}(d)^{2}),$
and hence
$E_{+}(d)-E_{-}(d)=-2\kappa_{0}\bigl(\delta_{+}(d)-\delta_{-}(d)\bigr)+o(e^{-\kappa_{0}d})=-4\kappa_{0}C\,e^{-\kappa_{0}d}+o(e^{-\kappa_{0}d}),$
which is equivalent to (7.4). ∎
Remark 7 (Agmon distance heuristic).
It is instructive to compare the exponential factor in (7.4) with the standard tunneling heuristics based on Agmon metrics. For a Schrödinger operator $\Delta+V(x)$ and an energy $E<\inf V$, the Agmon distance between two points $x$ and $y$ is defined by
$d_{E}(x,y)=\inf_{\gamma}\int_{0}^{1}\sqrt{(V(\gamma(t))-E)_{+}}\,|\dot{\gamma}(t)|\,dt,\qquad(\cdot)_{+}=\max(\cdot,0).$
In the present model the potential vanishes in the gap region, so $V=0$ between the shells, while $E=-\kappa_{0}^{2}<0$. Hence $V-E=\kappa_{0}^{2}$ in the gap and the integrand becomes the constant $\sqrt{-E}=\kappa_{0}$. Therefore the Agmon action across $(R_{1},R_{2})$ equals
$d_{E}(R_{1},R_{2})=\int_{R_{1}}^{R_{2}}\kappa_{0}\,dr=\kappa_{0}d,$
and the splitting scale $\exp(-d_{E}(R_{1},R_{2}))=\exp(-\kappa_{0}d)$ agrees with the explicit asymptotics in Theorem 17. Here $\exp(-2d_{E})$ corresponds to a tunneling probability (a squared amplitude), while the energy splitting is proportional to an overlap amplitude and therefore has the scale $\exp(-d_{E})$.
Acknowledgements
The author is grateful to Professor Pavel Exner for drawing his attention to earlier results on concentric $\delta$–shell interactions obtained by J. Shabani and collaborators. This comment led to a substantial clarification of the historical context and to an improvement of the introduction.
Appendix
Appendix A Calibration with semiconductor quantum dots
This appendix provides a scale and trend check for the double $\delta$–shell model based on representative parameters for core–shell quantum dots. It is not intended as a quantitative fit to any specific sample. We focus on the conduction–band profile for the electron. Optical transition energies, however, depend on additional ingredients such as the valence–band structure, excitonic Coulomb attraction, and polarization or self–energy effects, which are not captured by the present model.
In realistic core–shell structures, the relevant confinement levels may in fact be quasi–bound once finite wells or barriers and further band–structure effects are taken into account. Accordingly, the comparison presented here is meant only to capture order–of–magnitude estimates and qualitative trends.
Interface modeling and relation to BDD. Band edges vary across a few atomic layers ($\sim 0.3\,\mathrm{nm}$), which is thin compared with nanocrystal length scales; see, e.g., [9, 20]. Hence we approximate a thin interfacial step by a surface $\delta$–interaction. In BDD–type effective mass models one imposes continuity of $\psi$ and of the flux $(1/m^{\ast})\partial_{n}\psi$; see, for example, [13]. Here we keep a constant $m^{*}$ in each calibration and absorb interfacial physics into an effective coupling $\alpha$; it should not be interpreted as a microscopic BDD parameter.
Figure 1: Typical Type I (CdSe/ZnS) band alignment.
Figure 2: Typical Type II (CdTe/CdSe) band alignment.
Figures 22 illustrate schematic Type I (CdSe/ZnS) and Type II (CdTe/CdSe) alignments; representative offsets can be found in [19] (see also [10] for typical core–shell systems). Approximating a thin step of height $\Delta V$ and width $w$ by $\Delta Vw\,\delta(|x|-R)$, we obtain the dimensionless strength
$\alpha\simeq\frac{\Delta V}{E_{0}}\,\frac{w}{L_{0}},\qquad E_{0}=\frac{\hbar^{2}}{2m^{*}L_{0}^{2}},\quad L_{0}=1\,\mathrm{nm}.$ (A.1)
Thus $\mathrm{sign}(\alpha)=\mathrm{sign}(\Delta V)$. For Type I we use $\alpha_{1}<0$ (attractive inner wall) and $\alpha_{2}>0$ (outer barrier), while for Type II we reverse the signs: $\alpha_{1}>0$ and $\alpha_{2}<0$.
Type I (CdSe/ZnS): strong confinement. Take $m^{*}\simeq 0.13m_{0}$ (CdSe), so $E_{0}\simeq 0.29\,\mathrm{eV}$ [20]. Choose a representative geometry $R_{1}=2.5\,\mathrm{nm}$, $d=1.0\,\mathrm{nm}$ ($R_{2}=3.5\,\mathrm{nm}$), consistent with typical experimental ranges [14, 21, 10]. Fix $\alpha_{1}<0$ so that the single–shell problem at $R_{1}$ supports an ss–wave bound state at a typical scale $\kappa\sim\pi/R_{1}$ (hence $|\alpha_{1}|=O(1)$). For the outer interface, (A.1) with a representative offset (e.g. $\Delta V\simeq 0.7\,\mathrm{eV}$ and $w\simeq 0.3\,\mathrm{nm}$) gives $\alpha_{2}=O(1)$. A representative set
$R_{1}=2.5,\quad R_{2}=3.5,\quad\alpha_{1}\simeq-2.5,\quad\alpha_{2}\simeq 0.7$
yields an ss–wave bound state with a confinement scale $\Delta E_{\mathrm{conf}}$ of order $10^{-1}\,\mathrm{eV}$ (in physical units), i.e. a strongly confined electron level. Optical shifts are reduced by Coulomb and polarization/self–energy corrections, each typically $\sim 0.1\,\mathrm{eV}$ for radii of a few nanometers [9], so a blue shift of order $10^{-1}\,\mathrm{eV}$ is plausible, without aiming at a quantitative fit.
Type II (CdTe/CdSe): shallow outer–shell state. Take $m^{*}\simeq 0.11m_{0}$ (CdTe/CdSe), so $E_{0}\simeq 0.35\,\mathrm{eV}$ [20]. With the same geometry and the Type II sign pattern, (A.1) with a representative downward offset (e.g. $\Delta V\simeq-0.8\,\mathrm{eV}$, $w\simeq 0.35\,\mathrm{nm}$) gives $\alpha_{2}=O(1)<0$, while a moderate inner barrier $\alpha_{1}=O(1)>0$ is natural. For instance,
$R_{1}=2.5,\quad R_{2}=3.5,\quad\alpha_{1}\simeq+2.5,\quad\alpha_{2}\simeq-0.8$
typically produces (when an ss–wave bound state exists) a very small $\kappa$ and hence a shallow level on the meV scale, localized mainly in the outer shell. This captures the qualitative Type II picture: weak electron confinement. Observed Type II optical red shifts (often a few $10^{-1}\,\mathrm{eV}$) depend crucially on bandgap differences and excitonic/polarization effects beyond the present effective surface–interaction model.
Conclusion. This calibration is only a magnitude/trend check: Type I yields a strongly confined ss–state, whereas the Type II sign pattern yields a shallow outer–shell state. Quantitative optical predictions require a multiband/exciton model and are beyond the scope of this paper.
Appendix B Explicit partial–wave matrices
In this appendix we record an explicit closed form of the partial–wave matrices $m_{\ell}(z)$ appearing in Lemma 9. The formula follows from the standard spherical–wave expansion of the free resolvent kernel.
Let $k=\sqrt{z}$ with $\text{{\rm Im}}k>0$. In the present paper we evaluate the free resolvent kernel $G_{z}$ only at points of the form $x=R_{i}\omega$ and $y=R_{j}\omega^{\prime}$ with $i,j\in\{1,2\}$ and $0<R_{1}<R_{2}$. In this case the radial factors depend only on $R_{\min(i,j)}$ and $R_{\max(i,j)}$, so the mixed terms with $i\neq j$ are symmetric in $(i,j)$. The corresponding spherical–wave expansion reads
$G_{z}(R_{i}\omega,R_{j}\omega^{\prime})=ik\sum_{\ell=0}^{\infty}\sum_{m=-\ell}^{\ell}j_{\ell}\bigl(kR_{\min(i,j)}\bigr)\,h^{(1)}_{\ell}\bigl(kR_{\max(i,j)}\bigr)\,Y_{\ell m}(\omega)\,\overline{Y_{\ell m}(\omega^{\prime})},$
for every $\omega,\omega^{\prime}\in S^{2}$, where $j_{\ell}$ and $h^{(1)}_{\ell}$ denote the spherical Bessel and Hankel functions, and $\{Y_{\ell m}\}$ are the standard spherical harmonics forming an orthonormal basis of $L^{2}(S^{2},d\omega)$.
Lemma 18.
Assume $0<R_{1}<R_{2}$. For $i,j\in\{1,2\}$ let $m_{ij}(z)$ be the single–layer operators with kernel $G_{z}(R_{i}\omega,R_{j}\omega^{\prime})$, that is,
$(m_{ij}(z)\varphi)(\omega)=\int_{S^{2}}G_{z}(R_{i}\omega,R_{j}\omega^{\prime})\,\varphi(\omega^{\prime})\,d\omega^{\prime}.$
Then each spherical harmonic $Y_{\ell m}$ is an eigenfunction of $m_{ij}(z)$. More precisely,
$m_{ij}(z)Y_{\ell m}=\mu^{(\ell)}_{ij}(z)\,Y_{\ell m},$
with
$\mu^{(\ell)}_{11}(z)=ik\,j_{\ell}(kR_{1})\,h^{(1)}_{\ell}(kR_{1}),\qquad\mu^{(\ell)}_{22}(z)=ik\,j_{\ell}(kR_{2})\,h^{(1)}_{\ell}(kR_{2}),$
and
$\mu^{(\ell)}_{12}(z)=\mu^{(\ell)}_{21}(z)=ik\,j_{\ell}(kR_{1})\,h^{(1)}_{\ell}(kR_{2}).$
Consequently, the $2\times 2$ block in Lemma 9 is
$m_{\ell}(z)=\left(\begin{array}[]{cc}ik\,j_{\ell}(kR_{1})\,h^{(1)}_{\ell}(kR_{1})&ik\,j_{\ell}(kR_{1})\,h^{(1)}_{\ell}(kR_{2})\\ ik\,j_{\ell}(kR_{1})\,h^{(1)}_{\ell}(kR_{2})&ik\,j_{\ell}(kR_{2})\,h^{(1)}_{\ell}(kR_{2})\end{array}\right).$